List of Participants

Aasma, Ants (Estonia) Abd Al-Kader, Gamal M. (Egypt) Abdel-Megied, Mohamed (Egypt) Abdounur, Oscar Joao (Brazil) Abe, Yoshihiro (Japan) Abel, Mati (Estonia) Abeles, Francine F. (USA) Abels, Herbert (Germany) Abert, Miklos (Hungary) Abkar, Ali (Sweden) Abramochkin, Eugeny G. (Russian

Federation) Abramov, Gueorgui V. (Germany) Abramovich, Dan (USA) Abreu, Miguel (Portugal) A’Campo-Neuen, Annette (Germany) Accascina, Giuseppe (Italy) Adelmann, Clemens (Germany) Adelson-Velsky, George M. (Israel) Agud, Lucia (Spain) Ahlswede, Rudolf (Germany) Ahluwalia, Daljit S. (USA) Ahuja, Om Parkash (Singapore) Aiena, Pietro (Italy) Aigner, Martin (Germany) Aigner, Mats (Sweden) Aikawa, Takanori (Japan) Aipanov, Shamsha A. (Kazakhstan) Aisagaliev, Serikbai A. (Kazakhstan) Ajtai, Miklos (USA) Akbari, Sajeed (Iran) Akinyele, Olusola (USA) Alain-Roger, Nkamnang (Germany) Alania, Levan A. (Russian Federation) Alaoui, Larbi (Morocco) Albers, Bettina (Germany) Albrecht, Raphael (Germany) Albu, Toma (USA) Alcantara-Bode, Julio (Peru) Aldous, David John (USA) Aleksandrov, Aleksandr G. (Russian

Federation)

Alesker, Semyon (Israel) Alessandrini, Lucia (Italy) Alestalo, Pekka (Finland) Alexeyev, Alexander A. (Russian

Federation) Aliev, Baimurod (Tajikistan) Allende Alonso, Sira Maria (Cuba) Alli, Vahid (Iran) Alonso-Tarrio, Leovigildo (Spain) Alpay, Safak (Turkey) Alpers, Burkhard (Germany) Alpers, Karsten (Germany) Alpert, Christian (Germany) Al-Rashed, Abdallah M. (Kingdom of

Saudi-Arabia) Alt, Franz L. (USA) Alt, Helmut (Germany) Al-Thukair, Fawzi A. (Kingdom of

Saudi-Arabia) Altmann, Doris (Germany) Altmann, Klaus (Germany) Aluffi, Paolo (USA) Alves, Gloria L. M. (Brazil) Alves, Manuel J. (Mozambique) Al-Zanaidi, Mansour Abdullah (Kuwait) Amir, Dan (Israel) Ammann, Bernd (Germany) Amosov, Grigori (Russian Federation) Ancona, Vincenzo (Italy) Anders, Igor (Ukraine) Anderson, Joel H. (USA) Anderson, Laura M. (USA) Ando, Kazutoshi (Germany) Ando, Shiro (Japan) Andreian Cazacu, Cabiria I. (Romania) Andrews, George Eyre (USA) Andrié, Manfred (Germany) Anichini, Giuseppe (Italy) Ansorge, Rainer (Germany) Antipin, Anatoly (Russian Federation) Aoki, Shigeru (Japan) Appell, Jürgen (Germany)

62 List of Participants

Appolinaire, Nzali (Germany) Araki, Huzihiro (Japan) Arancibia, Jacqueline (Brazil) Aravinda, Chalya S. (India) Arbarello, Enrico (Italy) Aripov, Mersaid (Uzbekistan) Arkin, Vadim I. (Russian Federation) Armas-Sanabria, Lorena (Mexico) Arnold, Hans-Joachim (Germany) Arnold, Ludwig (Germany) Arnold, Vladimir Igorevich (Russian

Federation) Arnoux, Pierre (France) Arredondo, Juan H. (Mexico) Arslanov, Marat-M. (Russian Federation) Artamonov, Vyacheslav A. (Russian

Federation) Artamonova, Irena I. (Russian Federation) Artemiadis, Nicolas (Greece) Arthur, James Greig (Canada) Artigue, Michele (France) Asada, Akira (Japan) Asada, Teruko (Japan) Asavanant, Jack (Thailand) Ascheuer, Norbert (Germany) Ash, J. Marshall (USA) Ashna, Amir Hossein (Iran) Ashour, Attia A. (Egypt) Asiedu-Addo, Samuel K. (Ghana) Askarkhani, Abdoreza (Iran) Askin, Suat (Russian Federation) Aslanyan, Arthur (Russian Federation) Aslund, Jan (Sweden) Asmuss, Svetlana (Latvia) Aspinwall, Paul S. (USA) Assiamoua, Kofi V. S. (Togo) Astala, Kari (Finland) Asymont, Inna M. (Russian Federation) Atela, Pau (USA) Aubert, Jean-Christophe (Switzerland) Aubert, Pierre, F. (France) Auer, Franz (Germany) Auer, Roland (Germany) Aulaskari, Rauno (Finland) Aung, Min (Thailand) Ausekle, Jelena (Estonia) Avdispahic, Muharem (Bosnia and

Herzegovina) Avellaneda, Marco (USA) Avritzer, Dan (Brazil)

Axt, Paul (USA) Ayoub, Christine Williams (USA) Ayoub, Raymond G. (USA) Ayupov, Shavkat A. (Uzbekistan) Azarpanah, Fariborz (Iran) Azevedo, Alberto C. P. (Brazil) Azimi, Parviz (Iran)

Baas, Nils A. (Norway) Babbitt, Donald G. (USA) Babenko, Vladyslav F. (Ukraine) Bach, Volker (Germany) Bachmann, Heidy (Germany) Bacopoulos, Alexis (Greece) Badiozzaman, Abdul J. (Iran) Baidyk, Tatyana N. (Ukraine) Baier, Georg (Germany) Baier, Stephan (Germany) Baillif, Mathieu (Switzerland) Baker, Matthew H. (USA) Bakonyi, Mihaly (USA) Bakushinskii, Anatolii Borisovich (Russian

Federation) Balaji, Vikraman (India) Balashevich, Nataliya V. (Belarus) Balashova, Galina S. (Russian Federation) Baldus, Frank (Germany) Balke, Ludwig (Germany) Ball, John M. (United Kingdom) Balogh, Jozsef (Hungary) Bamba, Siaka Kante (Ivory Coast) Banaszak, Grzegorz M. (Poland) Bancerek, Grzegorz (Poland) Banchoff, Thomas F. (USA) Bandelow, Christoph (Germany) Bandt, Christoph (Germany) Bangert, Victor (Germany) Bangyan, Wen (P. R. China) Bank, Bernd (Germany) Bank, Peter (Germany) Banthien, Alexander (Germany) Bantsur, Nataliya (Ukraine) Baouendi, Salah (USA) Barannyk, Lyudmyla L. (Ukraine) Baranov, Alexander (Belarus) Barberis, María L. (Argentina) Barbu, Anca Stefania (Romania) Barchini, Leticia I. (USA) Baribaud, Claire M. C. (Switzerland) Barinka, Arne (Germany)

Barner, Klaus (Germany) Barrett, David E. (USA) Barrow-Green, June E. (United Kingdom) Bartels, Stephan (Germany) Barthel, Gottfried (Germany) Bartkiewicz, Monika (Poland) Bartolini Bussi, Mariolina (Italy) Bartosiewicz, Zbigniew J. (Poland) Bartsch, Thomas (Germany) Bass, Hyman (USA) Batistela, Claudia Helena F. (Brazil) Batyrev, Victor Vadimovich (Germany) Baudisch, Andreas (Germany) Bauer, Joachim (Germany) Bauhardt, Wolfgang (Germany) Baum, Helga (Germany) Baumann, Astrid (Germany) Bavula, Vladimir (United Kingdom) Bayer, Tilman (Germany) Bayoumi, Aboubakr (Kingdom of

Saudi-Arabia) Bazarov, Serguei (Russian Federation) Becker, Hans-J. (Germany) Becker, Paul-Georg (Germany) Behforooz, Hossein (USA) Behncke, Horst (Germany) Behrend, Kai A. (Canada) Behrends, Ehrhard (Germany) Belabas, Karim (France) Belinsky, Eduard (Zimbabwe) Bellaiche, Andre (France) Belousov, Evgeny G. (Russian Federation) Ben Hariz, Samir (France) Benjamin, Elliot (USA) Ben Massaoud, Hedi (Tunesia) Bensic, Mirta (Croatia) Benyash-Krivetz, Valerii Vaclavovich

(Belarus) Ben-Zvi, David (USA) Berenstein, Carlos A. (USA) Berezovskaya, Faina S. (Russian

Federation) Berg, Christian (Denmark) Berger, Bonnie (USA) Bergmann, Wolfgang Reinhold (Italy) Bergweiler, Walter (Germany) Berkovich, Vladimir (Israel) Berkovits, Juha (Finland) Berninger, Heiko (Germany) Bernstein, Joseph (Israel)

Beschler, Edwin (USA) Bessenrodt, Christine (Germany) Best, Christoph (Germany) Bethuel, Fabrice (France) Beutelspacher, Albrecht F. (Germany) Beutelspacher, Christoph (Germany) Bey, Christian (Germany) Beylkin, Gregory (USA) Bezuglyi, Sergey (Ukraine) Bhat, B. V. Rajarama (India) Bhattacharyya, Prodipeswar (India) Bhosle, Ushadevi N. (India) Bhupal, Mohan (Germany) Bierstedt, Klaus D. (Germany) Billard, Lynne (USA) Biller, Harald (Germany) Binder, Christa (Austria) Binder, Ilia A. (USA) Binder, Thomas S. (Germany) Birnir, Bjorn (USA) Bismut, Jean-Michel (France) Bissantz, Nicolai (Germany) Biswas, Indranil (India) Biucchi, Gabriele (Italy) Bjelica, Momcilo (Yugoslavia) Blackwood, Carol-Ann (USA) Blasius, Don (USA) Blatter, Christian (Switzerland) Blattner, Robert J. (USA) Bless, Michael H. (Germany) Bley, Andreas (Germany) Blickle, Manuel (USA) Blinov, Alexei V. (Russian Federation) Bloch, Spencer J. (USA) Blömker, Dirk (Germany) Boalch, Philip P. (United Kingdom) Bobenko, Alexander I. (Germany) Boche, Holger (Germany) Böcherer, Siegfried (Germany) Bock, Hans (Germany) Bock, Hans Georg (Germany) Böckle, Gebhard (Germany) Boden, Hans U. (USA) Bodfish, Ed (USA) Bödigheimer, Carl-Friedrich (Germany) Bodnarescu, M. V. (Germany) Boese, Fritz (Germany) Böge, Sigrid (Germany) Bognar, Gabriella (Hungary) Bogomolny, Eugene (France)

Bogovskii, Mikhail E. (Russian

Federation) Bohle, Christoph (Germany) Bojarski, Bogdan (Poland) Bolger, Robert E. (USA) Bolibok, Krzysztof M. (Poland) Bollobás, Béla (United Kingdom) Bonckaert, Patrick (Belgium) Bonet, Jose (Spain) Bonk, Mario (Germany) Bonnet, Philippe (France) Boo, Per-Anders (Sweden) Booß-Barnbek, Bernhelm (Denmark) Borchardt, Juergen (Germany) Borcherds, Richard E. (United Kingdom) Bordag, Ljudmila A. (Germany) Börgens, Manfred (Germany) Börger, Reinhard (Germany) Borgs, Christian (USA) Borisovich, Andrei Yu. (Poland) Borisovich, Yuri G. (Russian Federation) Borndörfer, Ralf (Germany) Borodin, Andrei Nikolaevich (Russian

Federation) Bors, Dorota (Poland) Borwein, David (Canada) Boskamp, Tobias (Germany) Bottazzini, Umberto (Italy) Böttcher, Roger (Germany) Boucherif, Abdelkader (Algeria) Bouchiba, Samir (Morocco) Bouetou, Thomas (Cameroon) Bouguima, Sidi Mohammed (Algeria) Boukricha, Abderrahman (Tunesia) Bourguignon, Jean-Pierre (France) Bourman, Iouri M. (Russian Federation) Boutot, Jean-Francois (France) Bovier, Anton (Germany) Boyallian, Carina (Argentina) Bozeman, James R. (USA) Bozer, Mehmet (Cyprus) Bozin, Vladimir (Yugoslavia) Braendli, Emil R. (Switzerland) Bramson, Maury D. (USA) Brandao, Paulo (Brazil) Brandenburg, Harald (Germany) Brandt, Manfred (Germany) Brandt, Stephan (Germany) Branner, Bodil (Denmark) Bratzler, Clemens (Germany)

Brauer, George (USA) Braun, Rüdiger W. (Germany) Braunß, Hans-Andreas (Germany) Brecht, Gerhard (Germany) Breckner, Wolfgang W. (Romania) Bredimas, Antoine-A. P. (France) Breger, Manfred (Germany) Brehm, Ulrich (Germany) Bremer, Ingo (Germany) Breuning, Manuel (Germany) Brezhnev, Yurii (Russian Federation) Bridges, Douglas S. (New Zealand) Brieden, Andreas (Germany) Brieskorn, Egbert (Germany) Brinzanescu, Vasile P. (Romania) Brodsky, Mikhail (USA) Brokate, Martin (Germany) Brookfield, Gary J. (USA) Broser, Britta (Germany) Brothers, Malgorzata (USA) Broughan, Kevin A. (New Zealand) Browder, Felix E. (USA) Brown, Aldric L. (India) Brown, Leon (USA) Brown, Tom C. (Canada) Brück, Rainer (Germany) Brucker, David (Israel) Bruckner, Gottfried (Germany) Brueckmann, Klaus Peter (Germany) Brunaud, Marc (France) Brüning, Jochen (Germany) Brunner, Götz (Germany) Brunotte, Horst (Germany) Bruns, Annika (Germany) Bruns, Peter (Germany) Bruns, Winfried (Germany) Brykalov, Sergei A. (Russian Federation) Brzezinski, Juliusz (Sweden) Bu, Charles (USA) Buan, Aslak B. (Norway) Bănulescu, C. Martha (Romania) Bubner, Nikolaus (Germany) Buchanan, Thomas (Germany) Buchholz, Detlev (Germany) Buchsteiner, Alexandra (Germany) Buchsteiner-Kießling, Edeltraud

(Germany) Buckwar, Evelyn (Germany) Buescu, Jorge C. (Portugal) Bujalance, Emilio (Spain)

Bujalance, Jose A. (Spain) Bulboaca, Teodor (Romania) Bulckens, Anne M. (Australia) Bulinski, Alexander V. (Russian

Federation) Bulinskij, Andrej V. (Russian Federation) Bumby, Richard T. (USA) Bungartz, Paul (Germany) Bunke, Olaf (Germany) Bunke, Ulrich (Germany) Burago, Dmitri (USA) Burde, Dietrich (Germany) Burger, Heinz H. (Germany) Burger, Isabella C. (South Africa) Burns, Daniel M. (USA) Busneag, Dumitru (Romania) Buswell, Stephen (United Kingdom) Butkovic, Davor (Croatia) Buttler, Michael (USA) Button, Jack O. (United Kingdom) Byrne, Catriona M. (Germany) Bytsko, Andrei G. (Russian Federation) Byunghan, Kim (USA)

Cabana, Alejandra (Venezuela) Cafarov, Dzhuma (Tajikistan) Cagliero, Leandro R. (Argentina) Cahen, Paul-Jean (France) Calkin, Neil J. (USA) Calvo, Jorge Albert (USA) Campbell, Paul J. (USA) Campillo, Antonio (Spain) Can, Mehmet (Turkey) Canales, Mónica del Pilar (Chile) Canas da Silva, Ana (USA) Cannon, Anne (Australia) Cannon, John J. (Australia) Cappel, Sylvan (USA) Caps, Oliver (Germany) Caraman, Petru (Romania) Cardiner, Anthony (United Kingdom) Cardona, Augusto Vieira (Brazil) Carlson, Robert (USA) Carmona, Angeles (Spain) Carnal, Henri C. (Switzerland) Caro, Jaime D. L. (Philippines) Carr, Jack (United Kingdom) Carroll, Robert W. (USA) Carter, J. Scott (USA) Casacuberta, Carles (Spain) Casati, Paolo (Italy)

Cascon, Ana (United Kingdom)

Case, Bettye A. (USA) Cassaigne, Julien (France) Castellet, Manuel (Spain) Catalan, Raquel-Garcia (Spain) Catanese, Fabrizio M. E. (Germany) Catinas, Emil A. (Romania) Cazaran, Jilyana (Australia) Cercignani, Carlo (Italy) Chadan, John (USA) Chahal, Jasbir S. (USA) Chaleyat-Maurel, Mireille (France) Chan, Tony F. (USA) Chang, Kun Soo (Korea) Chari, Manoj (Germany) Chatterji, Srishti (S. D.) (Switzerland) Chatyrko, Vitali (Sweden) Chayes, Jennifer T. (USA) Chebanov, Dmitriy A. (Ukraine) Chebotarev, Alexander M. (Russian

Federation) Chechkin, Gregory A. (Russian

Federation) Cheddadi, Abdelkhalek (Morocco) Chemla, Karine (France) Chen, Chin-Yun (Germany) Chen, Jiecheng (P. R. China) Chen, Wenxiong (USA) Chen, Zhiming (P. R. China) Cheng, Chong-Qing (P. R. China) Cheng, Ningning (P. R. China) Chentsov, Alexander G. (Russian

Federation) Cherednik, Ivan V. (USA) Cherinda, Marcos (Mozambique) Cherniha, Roman M. (Ukraine) Chernousov, Vladimir (Belarus) Cherry, William A. (Germany) Chiang, Tzuu-Shuh (Taiwan) Chien, Mao-Ting (Taiwan) Chikrii, Arkadii A. (Ukraine) Chillag, David (Israel) Chin, Angelina Y. M. C. (Malaysia) Chinen, Koji (Japan) Chiu, Sou-yung (Taiwan) Cho, Yong S. (Korea) Choban (Cioban), Mitrofan M. (Moldova) Chowdhury, Munibur Rahman

(Bangladesh) Christ, Michael (USA)

Christ, Ulrich (Germany) Christiansen, Chr. Andreas (Norway) Chua, Kok Seng (Singapore) Chynkulyak, Nataliya Nikolaevna

(Ukraine) Ciegis, Raimondas (Lithuania) Ciesielski, Zbigniew (Poland) Cieslik, Dietmar (Germany) Cioranescu, Doina (France) Cipu, Mihai I. (Romania) Ciric, Miroslav D. (Yugoslavia) Clarke, Francis W. (United Kingdom) Coates, John H. (United Kingdom) Cohen, Paula B. (France) Cohn, Harvey (USA) Colding, Tobias Holck (USA) Collet, Pierre (France) Colli, Eduardo (France) Colliot-Thélene, Jean-Louis (France) Colmez, Pierre (France) Colonius, Fritz (Germany) Conduché, Daniel (France) Conlon, Lawrence W. (USA) Connett, William C. (USA) Conrad, Marc (Germany) Consani, Katia (United Kingdom) Conte, Alberto (Italy) Cook, William J. (USA) Cooke, Roger L. (USA) Coonce, Harry B. (USA) Coornaert, Michel (France) Cordaro, Paulo (Brazil) Cordes, Frank (Germany) Cordes, Heinz O. (USA) Cornalba, Maurizio (Italy) Costa, Sueli I. R. (Brazil) Cowles, Lauren (USA) Crapo, Henry (France) Craveiro de Carvalho, Francisco J.

(Portugal) Crespo, Teresa (Spain) Crespo Crespo, Cecilia R. (Argentina) Cristea, Valentin Gabriel (Romania) Croft, Hallard T. (United Kingdom) Cromme, Ludwig (Germany) Cullen, Helen F. (USA) Cuntz, Joachim (Germany) Czarnecki, Maciej A. (Poland) Czichowski, Günter (Germany)

Dahan-Dalmedico, Amy (France)

Dais, Dimitrios I. (Germany) Dale, Knut Th. (Norway) Dalitz, Wolfgang (Germany) Damanik, David (Germany) Damlamian, Alain (France) Danchev, Daniel Marinov (Bulgaria) D’Ancona, Piero A. (Italy) Danet, Nicolae (Romania) Danet, Rodica Mihaela (Romania) Danishevs’kyy, Vladyslav V. (Ukraine) Darafsheh, Mohammad Reza (Iran) Dark, Rex (Ireland) Das, Laxminarayan (India) Dascalescu, Sorin (Romania) Dashkova, Ol’ga Yu. (Ukraine) Dasic, Vucic (Yugoslavia) Dauben, Joseph (USA) Daverman, Robert (USA) David, Herbert A. (USA) Davis, Chandler (Canada) Davis, James F. (USA) Davis, Philip J. (USA) Davydov, Alexei (Russian Federation) Dawson, Donald (Canada) Dawson, John W. (USA) Daxlberger, Christian (Germany) Debicka, Joanna (Poland) Debicki, Krzysztof (Poland) Debinska-Nagorska, Anna (Poland) De Blasi, Francesco (Italy) Debnath, Joyati (USA) Debnath, Narayan C. (USA) Dediu, Luminita Simona (New Zealand) de Faria, Edson (Brazil) de Guzman, Miguel (Germany) Deift, Percy Alec (USA) Deitmar, Anton (Germany) de Jong, Aise J. (USA) de Jong, Johan (Netherlands) Del Angel R., Pedro L. (Mexico) Deligne, Pierre (USA) Dell’ Antonio, Gianfausto (Italy) de Longueville, Mark (Germany) Del Riego, Lilia (Mexico) de Melo, Welington (Brazil) Demidenko, Gennadii V. (Russian

Federation) Demidov, Sergei S. (Russian Federation) Demuth, Michael (Germany) Deninger, Christopher (Germany)

Dennis, R. Keith (USA) Denzler, Jochen (Germany) De Sousa-Dias, M. Esmeralda (Portugal) Desquith, Etienne (Ivory Coast) Dessai, Anand (Germany) Deszcz, Ryszard (Poland) Deuber, Walter (Germany) Deuflhard, Peter (Germany) Deuschel, Jean-Dominique (Germany) Diaconis, Persi W. (USA) Diaconu, Calin A. (USA) Didenko, Victor D. (Ukraine) Diener, Karl-Heinz (Germany) Diep, Do Ngoc (Vietnam) Dieulefait, Luis V. (Spain) Dijkgraaf, Robbert (Netherlands) Dikoussar, Vassili (Russian Federation) Dillen, Franki (Belgium) Dimitric, Ivko M. (USA) Dinar, Nathan (Israel) Ding, Weiyue (P. R. China) Dinges, Hermann (Germany) Di Piazza, Luisa (Italy) Dittrich, Bianca (Germany) Djokovic, Dragomir (Canada) Dold, Albrecht (Germany) Dolzmann, Andreas (Germany) Domanski, Pawel (Poland) Domingo-Vecchioni, Miguel (Germany) Domokos, Matyas (Hungary) Donaldson, Simon K. (USA) Donig, Jörg (Germany) Dorn, Andreas (Germany) Dorninger, Dietmar (Austria) Dorschfeldt, Christoph (Germany) Doser, Jürgen (Germany) Dosiyev, Adigezal (Cyprus) Doslic, Tomislav (Croatia) Dotti, Isabel G. (Argentina) Doty, Stephen R. (USA) Douady, Adrien (France) Douady, Regine (France) Douglas, Lloyd (USA) Douglas, Robert J. (USA) Döweling, Sebastian (Germany) Downey, Jessica (USA) Doyen, Jean V. (Belgium) Drabant, Bernhard (United Kingdom) Draexler, Peter (Germany) Dragovic, Vladimir (Yugoslavia)

Dranishnikov, Alexander N. (USA)

Draper, Cristina (Spain)

Dress, Andreas W. M. (Germany) Dreyer, Wolfgang (Germany) Driver, Kathy A. (South Africa) Droescher, Christian (Germany) Druetta, Maria Josefina (Argentina) Drygas, Hilmar (Germany) Dryuma, Valery (Moldova) Duan, Jinqiao (USA) Dubickas, Arturas (Lithuania) Dubinsky, Andrej Yu. (Russian

Federation) Dubischar, Daniel (Germany) Dubrovin, Boris (Italy) Dudley, Richard M. (USA) Dudnikova, Tatjana (Russian Federation) Duduchava, Roland (Georgia) Duistermaat, Johannes J. (Netherlands) Duke, William (USA) Dümmel, Fritz (Germany) Dumortier, Freddy (Belgium) Dunne, Edward G. (USA) Duplij, Steven (Ukraine) Duren, Peter (USA) Durfee, Alan H. (USA) Dusembaev, Anuar (Kazakhstan) Duzaar, Frank (Germany) Dvoretzky, Aryeh (Israel) Dvornicich, Roberto (Italy) Dvorsky, Alexander (USA) Dwyer, William Gerard (USA) Dymara, Jan (Poland) Dynkin, Evsei M. (Israel) Dzhumadil’daev, Askar (Kazakhstan) Dzhuraev, Abduhamid (Tajikistan) Dziri, Raja (Tunesia) Dziubanski, Jacek (Poland)

Eberhard, Walter (Germany) Eberlein, Dominik (Germany) Eberlein, Ernst W. (Germany) Ebihara, Madoka (Japan) Ebmeyer, Carsten (Germany) Eckner, Stephan (Germany) Eckstein, Frank (Germany) Edwards, Elisabeth J. (United Kingdom) Edwards, Robert D. (USA) Eglewski, Dennis W. (USA) Egorov, Alexandre A. (Russian

Federation)

Eguchi, Kazuo (Japan) Ehret, Marietta (Germany) Ehrig, Hartmut (Germany) Ehrig, Rainald (Germany) Eibeck, Andreas (Germany) Eida, Atsuhiko (Japan) Eilertsen, Stefan (Sweden) Eisenblätter, Andreas (Germany) Eisenbrand, Friedrich (Germany) Eisenbud, David (USA) Eisentraeger, Kirsten (USA) Eisermann, Volker (Germany) Ekama Sossa Mben, Andre Sam

(Germany) Ekeland, Ivar (France) Ekhaguere, Godwin (Nigeria) Ekong, Samuel (France) Elbaz-Vincent, Philippe (France) Eliashberg, Yakov (USA) Eliasson, Håkan (Sweden) El Kinani, El Hassan (Morocco) Elliott, George A. (Canada) Elschner, Johannes (Germany) Elstrodt, Jürgen (Germany) El Yacoubi, Nouzha (Morocco) Emerson, Annette (USA) Emmer, Michele (Italy) Encinas, Andres, M. (Spain) Engel, Konrad (Germany) Engell, Sebastian (Germany) Engler, Hans (USA) Engquist, Bjorn Erik (Sweden) Enkhbat, Rentsen (Mongolia) Enock, Michel (France) Enomoto, Kazuyuki (Japan) Enß, Volker (Germany) Epstein, Mordechai (Israel) Erbay, Hüsnü Ata (Turkey) Erbay, Saadet (Turkey) Erdmann, Bodo (Germany) Erdogan, Hakki Ismail (Turkey) Eriksson, Folke S. (Sweden) Erle, Dieter (Germany) Eschenbach, Christina (Germany) Eschenburg, Rolf (Germany) Eskin, Alexander (USA) Esser, Angelika E. (Germany) Esteves, Eduardo S. (Brazil) Estevez, José Luis (Spain) Estevez Schwarz, Diana (Germany)

Eto, Kazufumi (Japan) Eudave-Munoz, Mario (Mexico) Evans, David E. (United Kingdom) Ewing, John H. (USA) Eyles, Joseph W. (USA) Ezquerro, Jose A. (Spain)

Faatz, Vanessa (Germany) Faber, Carel (USA) Fabian, Benjamin (Germany) Faddeev, Ludvig D. (Russian Federation) Fahl, Michael (Germany) Fahsi, Abdelhak (Morocco) Faltings, Gerd (Germany) Faminskii, Andrei V. (Russian Federation) Fandom Noubiap, Roger (Germany) Farge, Marie (France) Farr, Graham E. (Australia) Farsi, Carla E. (USA) Fasso Velenik, Agnese (Germany) Fathi, Albert (France) Fayad, Bassam (France) Faybusovich, Leonid (USA) Fedosov, Boris (Germany) Fedotov, Igor (South Africa) Fedotov, Sergei (United Kingdom) Fedynyak, Stepan (Ukraine) Feehan, Paul M. N. (USA) Feher, Franziska (Germany) Feichtner, Eva Maria (Germany) Feigenbaum, Joan (USA) Feith, Michael (Germany) Feix, Birte (United Kingdom) Feldhusen, Dirk (Germany) Fels, Marten (Germany) Fel’shtyn, Alexander (Germany) Felsner, Stefan (Germany) Fernandes, Cristina G. (Brazil) Fernandes, Rui L. (Portugal) Fernandez, Arturo (Spain) Feroe, John A. (USA) Ferreira, Fernando Flavio (Portugal) Ferreira, J. A. (Portugal) Ferus, Dirk (Germany) Fialowski, Alice (Hungary) Fiedler, Bernold (Germany) Fieker, Claus (Germany) Fieseler, Martin (Germany) Figiel, Tadeusz (Poland) Filip, Anja (Germany) Filipsson, Lars (Sweden)

Fillmore, Peter A. (Canada) Fimmel, Torsten (Germany) Fintushel, Ronald A. (USA) Fischer, Alexander (Germany) Fischer, Andreas (Germany) Fischer, Arthur E. (USA) Fischer, Gerd (Germany) Fishel, Susanna D. (Germany) Fittler, Robert (Germany) Fjelstad, Paul (USA) Flach, Matthias (USA) Flato, Moshé (France) Fleischmann, Klaus (Germany) Fleischmann, Peter (Germany) Flor, Peter (Austria) Flores-Bazan, Fabian (Chile) Florit, Luis A. (Brazil) Flytzanis, Elias (Greece) Folland, Gerald B. (USA) Föllmer, Hans (Germany) Fontana, Marco (Italy) Foreman, Matthew (USA) Forger, Michael (Brazil) Förster, Christian (Germany) Förster, Karl-Heinz (Germany) Forti, Marco (Italy) Fossi Talom, Leopold (Germany) Fossum, Robert M. (USA) Fotheringham, Gerhard (Germany) Fourie, Jan H. (South Africa) Francis, George (USA) Frank, András (Hungary) Frank, Michael (USA) Franke, Dirk (Germany) Frankl, Astrid (Germany) Franz, Matthias (Germany) Freedman, Michael Hartley (USA) Fregier, Yael (France) Freidlin, Mark (USA) Frese, Hans-Hermann (Germany) Frey, Christian (Germany) Friedland, Lew (USA) Friedlander, Eric Mark (USA) Friedlander, Susan (USA) Friedman, Eduardo C. (Chile) Friedrich, Axel (Germany) Friedrichs, Carsten (Germany) Friese, Nicolas (Germany) Friese, Stephan (Germany) Fring, Andreas (Germany)

Fritsch, F. Rudolf (Germany) Fritzlar, Torsten (Germany) Frolovitchev, Sergey (Russian Federation) Fry, Robb C. (Canada) Frydrych, Mariusz (Poland) Fuchs, Ulrich (Germany) Fuhrmann, Jürgen (Germany) Fujii, Akio (Japan) Fujii, Nobuo (Japan) Fujita, Keiko (Japan) Fukushima, Masatoshi (Japan) Fukushima, Nobuhisa (Japan) Furihata, Daisuke (Japan) Furutani, Kenro (Japan)

Gaál, István (Hungary) Gaertner, Klaus (Germany) Gaiko, Valery A. (Belarus) Gajda, Wojciech J. (Poland) Gajewski, Herbert (Germany) Galewski, Marek (Poland) Gallavotti, Giovanni (Italy) Gallot, Sylvestre (France) Gamba, Irene M. (USA) Gamst, Jens (Germany) Gan, Shaobo (P. R. China) Gangl, Herbert (Germany) Ganief, M. Shahiem (South Africa) Gao, Xuhong (USA) Garimella, Venkatalakshmi Gayatri

(India) Garling, David J. H. (United Kingdom) Garrido Bullon, Angel L. (Spain) Gärtner, Bernd (Switzerland) Gärtner, Jürgen (Germany) Gatermann, Karin (Germany) Gattazzo, Remo (Italy) Gatto, A. Eduardo (USA) Gawrilow, Ewgenij (Germany) Gay, David M. (USA) Gebauer, Susanna (Germany) Gebel, Michael (Germany) Geddes, Keith O. (Canada) Gehring, Frederick W. (USA) Geiger, Jochen (Germany) Geisser, Thomas H. (Germany) Geissler, Katharina (Germany) Gelfand, Sergei (USA) Geller, Daryl N. (USA) Geluk, Jaap (United Arab Emirates) Georgescu, Adelina (Romania)

Georgii, Hans-Otto (Germany) Gérard, Olivier P. (France) Gerisch, Magnus (Germany) Gerlach, Eberhard (USA) Gerstner, Manfred (Germany) Geyler, Vladimir A. (Russian Federation) Ghanaat, Patrick (Germany) Ghate, Eknath P. (India) Ghazel, Moncef (Tunesia) Ghorpade, Sudhir R. (India) Ghosh, Jayanta Kumar (India) Ghoussoub, Nassif (Canada) Giaquinto, Anthony (USA) Gidea, Marian (USA) Giesen, Joachim (Switzerland) Gietl, Thomas (Germany) Gigena, Salvador D. R. (Argentina) Gil, Juan B. (Germany) Gilanyi, Attila (Hungary) Gilliard, Elizabeth (United Kingdom) Gillman, Leonard (USA) Gillot, Antonio E. (Guatemala) Gil-Medrano, Olga (Spain) Giordano, Anna (Italy) Giorgilli, Antonio (Italy) Giudici, Reinaldo (Venezuela) Gjone, Gunnar (Norway) Glass, Thomas (Germany) Glazman, Mary (Mexico) Gleska, Alina (Poland) Glimm, Ekkehard (Germany) Glitzky, Annegret (Germany) Glöckner, Helge (Germany) Gloden, Raoul R. F. G. (Luxembourg) Glutsuk, Alexei A. (Russian Federation) Göbel, Dietmar (Germany) Göbel, Silke (Germany) Goddard, Peter (United Kingdom) Goemans, Michel X. (Belgium) Goette, Sebastian (Germany) Goetz, Abraham (USA) Goldblatt, Robert (New Zealand) Goldstern, Martin (Austria) Gollek, Hubert (Germany) Golodets, Valentyn Ja. (Ukraine) Golomb, Michael (USA) Goloubeva, Valentina (Russian

Federation) Golovanov, Anton A. (Russian Federation) Golovanova, Nina F. (Russian Federation)

Gomenyuk, Sergey (Ukraine) Gomez Bermudez, Carlos (Spain) Gomez Bofill, Walter (Germany) Gomez-Mont, Xavier (Mexico) Gomez Pardo, José L. (Spain) Gonchar, Andrei A. (Russian Federation) González, Cesareo (Spain) Goodman, Gerald S. (Italy) Goodman, Jacob E. (USA) Gorenflo, Rudolf (Germany) Goresky, R. Mark (USA) Goritsky, Andrey Yu. (Russian

Federation) Gorodski, Claudio (Brazil) Gottlieb, Daniel H. (USA) Gottschalk, Harald (Germany) Götze, Friedrich (Germany) Gowers, W. Timothy (United Kingdom) Grabitz, Martin (Germany) Graeff, Robert (Germany) Graf, Gian Michele (Switzerland) Graf, Klaus-D. (Germany) Graham, Ronald L. (USA) Gramain, François (France) Gramsch, Bernhard F. (Germany) Granovsky, Boris L. (Israel) Grase, Saida (Germany) Grassi, Michele (Italy) Grassmann, Hubert (Germany) Gräter, Joachim (Germany) Grattan-Guiness, Ivor (United Kingdom) Gray, Jeremy John (United Kingdom) Grayson, Daniel R. (USA) Green, Mark L. (USA) Greengard, Leslie Frederick (USA) Greenspoon, Arthur (USA) Greil, Anton (Germany) Greither, Cornelius (Canada) Grell, Juliana (Germany) Grenander, Ulf (USA) Greuel, Bernd (Germany) Greuel, Gert-Martin W. (Germany) Grieser, Daniel (Germany) Griffiths, Peter L. (United Kingdom) Grinberg, Natalia I. (Germany) Grinnell, Raymond J. (Barbados) Gritzmann, Peter (Germany) Grkovska, Slavica (Macedonia) Grobler, Jacobus J. (South Africa) Grobstich, Peter (Germany)

Gröwe-Kuska, Nicole (Germany) Grof, Jozsef (Hungary) Gröger, Konrad (Germany) Grölz, Wolfgang (Germany) Gromov, Nikolai (Russian Federation) Gronau, Hans-Dietrich O. F. (Germany) Gröpl, Clemens (Germany) Gropp, Harald (Germany) Grosse-Erdmann, Karl-Goswin (Germany) Grotowski, Joseph F. (Germany) Grötschel, Martin (Germany) Gruber, Michael J. (Germany) Grund, Friedrich (Germany) Grundhöfer, Theo (Germany) Grünvogel, Stefan Michael (Germany) Grüter, Michael (Germany) Grzanna, Jürgen (Germany) Grzaslewicz, Ryszard (Poland) Grzymkowski, Radoslaw (Poland) Gu, Weiqing (USA) Guccione, Linda (USA) Guddat, Jürgen (Germany) Guillaume, Anne (United Kingdom) Guillopé, Laurent (France) Gunawan, Hendra (Indonesia) Gundlach, Karl-Bernhard (Germany) Gunson, Jack (United Kingdom) Günzler, Hans (Germany) Guo, Jin Yun (P. R. China) Guo, Shirong (United Kingdom) Gurak, Stanley J. (USA) Gursky, Matthew J. (USA) Gusein-Zade, Sabir M. (Russian

Federation) Gusevskii, Nikolai (Brazil) Gusinde, Ellinor (Germany) Gustedt, Jens (Germany) Guthschmidt, Norman (Germany) Gutiérrez, José M. (Spain) Guyard, Frederic (Germany)

Haase, Christian (Germany) Haase, Markus (Germany) Habetha, Klaus (Germany) Haccou, Maryene (Netherlands) Hachimori, Masahiro (Japan) Hackbusch, Wolfgang (Germany) Hafner, Paul R. (New Zealand) Hag, Kari (Norway) Hag, Per (Norway) Haida, Minoru (Japan)

Haidar, Mahmoud (Kuwait) Haines, Thomas J. (Germany) Hajdu, Lajos (Hungary) Hajduk, Boguslaw (Poland) Hajnal, Peter (Hungary) Hajto, Zbigniew (Spain) Hakim, Jeffrey L. (USA) Hakulinen, Ville (Finland) Hales, Thomas C. (USA) Hall, Peter Gavin (Australia) Hall, Richard L. (Canada) Haller, Rainis (Estonia) Halperin, Benjamin D. (USA) Halupczok, Karin (Germany) Hamana, Yuji (Japan) Hämarik, Uno (Estonia) Han, Chong-Kyu (Korea) Han, Zheng-Chao (USA) Handa, Kenji (Japan) Handrock-Meyer, Sybille (Germany) Hanke, Bernhard (Germany) Hanke, Timo (Germany) Hannappel, Susanne (Germany) Hansen, Sönke (Germany) Hansen, Vagn L. (Denmark) Hansen, Wolfhard (Germany) Hantschmann, Andreas (Germany) Happel, Dieter (Germany) Harbater, David (USA) Harmand, Peter (Germany) Harms, Jan (Germany) Haroutunian, Samvel (Armenia) Harris, Michael H. (France) Härterich, Jörg (Germany) Härterich, Martin (Germany) Hartl, Urs T. (Germany) Hartmann, Stephan (Germany) Hartmann, Sven (Germany) Hartz, David (USA) Harutjunjan, Gohar (Germany) Harzheim, Egbert R. (Germany) Hasegawa, Keizo (Japan) Hasegawa, Kenji (Japan) Hasenjaeger, Gisbert F. R. (Germany) Hashagen, Ulf (Germany) Hashiguchi, Hideko (Japan) Hashiguchi, Norikazu (Japan) Hashimoto, Hideya (Japan) Hass, Joel (USA) Hassler, Uwe (Germany)

Håstad, Johan (Sweden) Hatori, Asako (Japan) Hattori, Akio (Japan) Haus, Utz-Uwe (Germany) Hausen, Jürgen (Germany) Hauser, Nico (Germany) Hausmann, Jean-Claude R. (Switzerland) Haverhals, John S. (USA) Hayashi, Shuhei (Japan) Hayes, Brian T. (USA) Hazewinkel, Michiel (Netherlands) He, Xinyu (United Kingdom) Hebermehl, Georg (Germany) Hebert, Michel (Egypt) Hedberg, Lars Inge (Sweden) Hedenmalm, Haakan P. (Sweden) Hege, Hans-Christian (Germany) Heiermann, Volker J. (Germany) Heiming, Helmut (Germany) Hein, Georg (Germany) Heinig, Hans P. (Canada) Heintze, Ernst (Germany) Heinz, Alois (Germany) Heinze, Joachim (Germany) Heiss, Stefan (Germany) Hejcman, Jan (Czech Republic) Heldermann, Norbert (Germany) Hélein, Frédéric (France) Helemskii, Alexander Ya. (Russian

Federation) Hell, Stephan (Germany) Helmberg, Christoph (Germany) Helmberg, Gilbert (Austria) Helmbold, Matthias (Germany) Helversen-Pasotto, Anna (France) Hemkemeier, Boris (Germany) Henk, Martin (Germany) Hennig, Franziska (Germany) Hennig, Steffen (Germany) Henrion, René (Germany) Heringlehner, Josef (Germany) Herman, Michael (France) Hermann, Frieder (Germany) Herrmann, Burghard F. G. (Germany) Herrmann, Eva (Germany) Herrmann, Oliver J. (Germany) Herrmann, Siegfried (Germany) Hertling, Claus (Germany) Hertrich-Jeromin, Udo J. (Germany) Heß, Florian (Germany)

Heuer, Gerald A. (USA) Heyer, Herbert K. W. (Germany) Hidaka, Fumio (Japan) Hidalgo, Ruben A. (Chile) Higson, Nigel David (USA) Hildebrand, Roland (Germany) Hille, Lutz (Germany) Hillman, Jonathan A. (Australia) Hilsberg, Isabel (Germany) Himonas, Alexandrou A. (USA) Hinrichs, Aicke (Germany) Hintermann, Thomas (Switzerland) Hinz, Andreas M. (Germany) Hinze, Annika (Germany) Hinze, Michael (Germany) Hirayama, Hiroshi (Japan) Hironaka, Eriko (USA) Hirzebruch, Friedrich E. P. (Germany) Hjorth, Gregory (USA) Hoang, Viet Ha (United Kingdom) Höbel, Willi (Germany) Hodgson, Bernard R. (Germany) Hofer, Helmut H. W. (Switzerland) Hoffmann, Detlev W. (France) Hoffmann, Karl-Heinz (Germany) Hoffmann, Sigrid (Germany) Hoffmann, Tim (Germany) Hoffmann, Werner (USA) Hofmann, Karl-Heinrich (Germany) Hofmann, Wolf (Germany) Hohberger, Horst H. (Germany) Hohmuth, Lars (USA) Höhn, Gerald (Germany) Holm, Thorsten (Germany) Holmann, Harald R. A. (Switzerland) Holme, Audun (Norway) Holopainen, Ilkka (Finland) Holst, Stefan (Germany) Holt, Fred B. (USA) Holte, John M. (USA) Holtkamp, Ralf (Germany) Holtmanns, Silke (Germany) Holz, Martin (Germany) Holzapfel, Rolf-Peter (Germany) Hömberg, Dietmar (Germany) Homburg, Ale Jan (Germany) Hooke, Nigel (Australia) Hoppensteadt, Frank (USA) Hora, Akihito (Japan) Horiuchi, Kiyomitsu (Japan)

Horiuchi, Ryutaro (Japan) Horiuchi, Toshio (Japan) Horn, Dietmar (Germany) Hornbostel, Jens (France) Hornor, William E. (USA) Horst, Ulrich (Germany) Horstmann, Dirk (Germany) Hotje, Herbert (Germany) Hou, Thomas Yizhao (USA) Hou, Zhanyuan (United Kingdom) Houh, Chorng Shi (USA) Houssni, Mohamed (Morocco) Hovhannisyan, Gro (Armenia) Hric, Roman (Slovakia) Hrushovski, Ehud (Israel) Hryniv, Ostap (Germany) Hu, Po (USA) Huang, Jing-Song (P. R. China) Huber, Arla M. (USA) Huebner, Friedrich-Karl

(Germany) Huebner, Marianne (USA) Hughes, Kenneth R. (South Africa) Huisinga, Wilhelm (Germany) Huisken, Barbara (Germany) Huisken, Gerhard (Germany) Huisman, Johannes (France) Hulek, Klaus W. (Germany) Hullet, Eduardo G. (Germany) Hünlich, Rolf (Germany) Hunt, John H. V. (South Africa) Hürlimann, Werner S. (Switzerland) Hurri-Syrjanen, Ritva (Finland) Hurtubise, Jacques (Canada) Husemoller, Dale (France) Hussain, Mansour G. (Kuwait) Hüsseinov, Farhad V. (Turkey) Hutchins, Carol (USA) Hutchinson, Kevin (Ireland) Hutchison, Jeanne S. (USA) Huth, Reinhardt (Germany) Hwang, Tea-Yuan (Taiwan)

Ibisch, Horst (France) Idczak, Dariusz (Poland) Igarashi, Masayuki (Japan) Ihara, Yasutaka (Japan) Il’ichev, Vitaly G. (Russian Federation) Ilolov, Mamadsho (Tajikistan) Imamoglu, Özlem (Switzerland) Im Hof, Hans-Christoph (Switzerland)

Imhof, Jean-Pierre (Switzerland) Imkeller, Peter (Germany) Inassaridze, Hvedri (Georgia) Ioffe, Dmitry (Germany) Ion, Patrick D. F. (USA) Ionescu, Cristodor P. (Romania) Iooss, Gerard M. (France) Iosifescu, Marius V. (Romania) Iranmanesh, Ali (Iran) Iricanin, Bratislav (Yugoslavia) Isbell, John (USA) Ischebeck, Friedrich (Germany) Ishigami, Yoshiyasu (Japan) Ishikawa, Takeshi (Japan) Ishiwata, Emiko (Japan) Ishizaki, Katsuya (Japan) Israel, Giorgio (Italy) Ito, Hiroyuki (USA) Ito, Ryuichi (Japan) Ito, Yukari (Japan) Itoh, Jin-ichi (Japan) Itoh, Tatsuo (Japan) Iturriaga, Renato (Mexico) Ivanov, Alexandr O. (Russian Federation) Ivanov, Sergei (USA) Ivansic, Ivan (Croatia) Ivarsson, Björn (Sweden) Iwasaki, Katsunori (Japan) Iwasaki, Yoshimitsu (Japan) Izadi, Elham (USA)

Jackson, Allyn (USA) Jacob, Niels (Germany) Jaeger, Arno (Germany) Jaehnisch, Michael (Germany) Jaffe, Arthur M. (USA) Jäger, Willi (Germany) Jagers, Albertus A. (Netherlands) Jahanshahi, Mohammad (Iran) Jahn, Olaf (Germany) Jahnke, Thomas (Germany) Jaiani, George (Georgia) Jakobsson, Stefan B. A. (Sweden) James, Donald G. (USA) James, Kevin (USA) Jang, Sophia R.-J. (USA) Janno, Jaan (Estonia) Jannsen, Uwe (Germany) Jansen, Karl-Heinz (Germany) Januszkiewicz, L. Tadeusz (Poland) Janz, Raimund (Switzerland)

Jarnik, Jiri (Czech Republic) Jarvis, Tyler J. (USA) Jean, Roger V. (Canada) Jeltsch, Rolf (Switzerland) Jensen, Christian U. (Denmark) Jensen, Robert R. (USA) Jensen, Ronald B. (Germany) Jeremias-Lopez, Ana (Spain) Jeschke, Sabina (Germany) Jewett, Rumara (USA) Ji, Lizhen (USA) Jiang, Renfang (USA) Jiang, Song (P. R. China) Jimenez, Miguel A. (Mexico) Jing, Naihuan (USA) Jochmann, Frank (Germany) Johansson, Bo I. (Sweden) John, Daniel (Germany) Johnson, Harold H. (USA) Johnson, Roy A. (USA) Johnson, William B. (USA) Johnston, Elizabeth A. (United Kingdom) Johnstone, Iain Murray (USA) Jones, Mark C. W. (United Kingdom) Jones, Vaughan (USA) Jorge, Maria H. (USA) Jorgensen, Palle E. T. (USA) Joseph, Benjamin S. (USA) Jost, Michael (Germany) Joswig, Michael (Germany) Jovanovic, Bosko (Yugoslavia) Joyce, Dominic (United Kingdom) Józefiak, Tadeusz (USA) Juhnke, Friedrich (Germany) Jun, Sungtae (Korea) Junek, Heinz (Germany) Jung, Heinz Adolf, H. A. (Germany) Jünger, Michael (Germany) Jungnickel, Dieter (Germany) Jurkschat, Henrik (Germany) Just, Andrzej (Poland) Jutila, Matti (Finland)

Kabanov, Alexandre (USA) Kaczor, Wieslawa J. (Poland) Kade, Wolfgang R. (Germany) Kadeishvili, Tornike (Georgia) Kadelburg, Zoran (Yugoslavia) Kadets, Volodymyr M. (Ukraine) Kadison, Richard V. (USA) Kaenders, Rainer H. (Germany)

Kahle, Reinhard (Germany)

Kahn, Bruno (France) Kahn, Donald (USA) Kaibel, Volker (Germany) Kairies, Hans-Heinrich (Germany) Kaiser, Hans-Christoph (Germany) Kajikiya, Ryuji (Japan) Kakiuchi, Nabuhiko (Japan) Kalita, Evgeni A. (Ukraine) Kallel, Sadok (Canada) Kallies, Jürgen (Germany) Kaluch I., Yurij (Ukraine) Kalus, Norbert (Germany) Kamal, Ahmed A. M. (Egypt) Kambayashi, Tatsuji (Japan) Kamont, Anna (Poland) Kamps, Klaus Heiner (Germany) Kamvissis, Spyridon (France) Kaneko, Makoto (Japan) Kang, Ming-Chang (Taiwan) Kangro, Raul (Estonia) Kangro, Urve (Estonia) Kantor, William M. (USA) Kanzaki, Teruo T. K. (Japan) Kappler, Fritz Otto (Germany) Kappos, Efthimios (United Kingdom) Kapranov, Mikhail M. (USA) Kapur, Aruna (India) Karamzadeh, Omid-Ali (Iran) Karbe, Manfred (Germany) Karcher, Hermann (Germany) Karlof, John K. (USA) Karolinsky, Eugene A. (Ukraine) Karoubi, Max (France) Karoui, Abderrazek (Tunesia) Kasaev, Anzor D. (Russian Federation) Kascelan, Vidosava (Yugoslavia) Kashani, S. M. B. (Iran) Kashiwada, Toyoko (Japan) Kasse, Bettina (Germany) Kassel, Christian (France) Kasyanov, Victor (Russian Federation) Katayama, Soichiro (Japan) Katona, Gyula O. H. (Hungary) Katre, Shashikant (India) Kats, Isaak Yakovlevich (Russian

Federation) Katsnelson, Victor, E. (Israel) Katz, Sheldon H. (USA) Kauffmann, Andreas (Germany)

Kauffmann, Louis H. (USA) Kaufmann, Ralph (France) Kaup, Burchard (Switzerland) Kaup, Ludger (Germany) Kausz, Ivan (Germany) Kawamata, Yujiro (Japan) Kawamura, Kiko (Japan) Kawano, Nichiro (Japan) Kawohl, Bernd (Germany) Kayll, Mark (USA) Kayumov, Ilgiz G. (Russian Federation) Kazantsev, Ivan G. (Russian Federation) Kazunori, Kikuchi (Japan) Kearton, Cherry (United Kingdom) Kecs, Wilhelm (Romania) Keedwell, A. Donald (United Kingdom) Keener, Lee L. (Canada) Kegel, Otto H. (Germany) Kellendonk, Johannes (Germany) Keller, Gerhard (Germany) Kemperman, Johannes H. (USA) Kenmotsu, Katsuei (Japan) Kerkour, Ahmed (Morocco) Kern, Uwe (Germany) Kersten, Ina (Germany) Ketelhut, Gunnar (Germany) Khalil, Zohel S. (Canada) Khan, Nazir M. (Pakistan) Khare, Chandrashekhar B. (India) Khelladi, Abdelkader (Algeria) Khilchenko, Lyudmila Y. (USA) Khimshiashvili, George N. (Poland) Khitrik, Mikhail (Sweden) Khokhlov, Vladimir (Russian Federation) Khorrampanahi, Mehrdad (USA) Khosrovshahi, Gholamreza B. (Iran) Khots, Boris S. (USA) Khrennikov, Andrei (Sweden) Kiechle, Hubert (Germany) Kifer, Yuri (Israel) Kiguradze, Ivan (Georgia) Kiguradze, Tariel (Georgia) Kikuchi, Shigeki S. K. (Japan) Kikuchi, Shigetaka S. K. (Japan) Kilsch, Dieter (Germany) Kim, Dohan (Korea) Kim, Hoil (Korea) Kim, Sang Moon (Korea) Kimura, Noriaki (Japan) Kimura, Takashi (USA)

Kinnmark, Ingemar (USA)

Kirchberg, Eberhard (Germany) Kirchgässner, Klaus (Germany) Kirchgraber, Urs (Switzerland) Kirchhoff, Dennis (Germany) Kirchner, Thomas (Germany) Kirrinnis, Peter (Germany) Kisaka, Masashi, M. K. (Japan) Kiselman, Christer O. (Sweden) Kisil, Vladimir V. (Ukraine) Kislyakov, Sergey (Russian Federation) Kissin, Edward (United Kingdom) Kister, James M. (USA) Kister, Jane E. (USA) Kitano, Teruaki (Japan) Kiyek, Karlheinz (Germany) Klassert, Steffen (Germany) Klazar, Martin (Czech Republic) Klein, Henning (Germany) Klein, Olaf (Germany) Kleshchev, Alexander S. (USA) Klesov, Oleg I. (Ukraine) Klimek, Malgorzata (Poland) Klimenko, Stanislav V. (Germany) Klin, Mikhail H. (Israel) Klingenberg, Christian (Germany) Klopsch, Benjamin (United Kingdom) Klüners, Jürgen (Germany) Knauf, Andreas (Germany) Kneissler, Jan (Germany) Knight, Frank B. (USA) Knobloch, Eberhard H. (Germany) Knopov, Pavel S. (Ukraine) Knuetel, Claudia D. (Germany) Ko, Hyoung J. (Korea) Kobayashi, Keiko (Japan) Koch, Helmut (Germany) Koch, Herbert (Germany) Koch, Robert (Germany) Koch, Tino (Germany) Kochol, Martin (Slovakia) Kochubei, Anatoly N. (Ukraine) Koda, Takashi (Japan) Kodiyalam, Vijay (India) Koenig, Wolfgang D. (Germany) Koepke, Peter G. (Germany) Koerner, Friederike (Germany) Kogiso, Takeyoshi (Japan) Kohaupt, Ludwig (Germany) Kohayakawa, Yoshiharu (Brazil)

Köhler, Ekkehard (Germany) Köhne, Hartmut (Germany) Köhnen, Walter (Germany) Kohr, Gabriela (Romania) Kohr, Mirela (Romania) Kojok, Badrie (Ivory Coast) Kokilashvili, Vakhtang M. (Georgia) Kokubu, Hiroshi (Japan) Kolk, Johan A. (Netherlands) Kolodyazhniy, Volodymyr M. (Ukraine) Kolomiets, Yuriy V. (Ukraine) Koltchinskii, Vladimir I. (USA) Kolumban, Iosif (Romania) Komatsu, Hikosaburo (Japan) Kometch, Alexandre (Russian Federation) Kondratiev, Anatolii S. (Russian

Federation) Kondratiev, Vladimir A. (Russian

Federation) König, Gerhard (Germany) König, Hermann (Germany) Konno, Hitoshi (Japan) Konnov, Valery (Russian Federation) Konrad, Alois (Germany) Kontsevich, Maxim (France) Koornwinder, Tom H. (Netherlands) Kopanskii, Alexander Ya. (Moldova) Koppelberg, Sabine (Germany) Koprucki, Thomas (Germany) Kopylov, Iaroslav A. (Russian Federation) Koranyi, Adam (USA) Kordyukov, Yuri A. (Russian Federation) Korepanov, Igor (Russian Federation) Korey, Michael B. (Germany) Korneichuk, Nikoloj P. (Ukraine) Kornhuber, Ralf (Germany) Kortesi, Peter (Hungary) Koschorke, Ulrich (Germany) Koshevoy, Gleb A. (Russian Federation) Koshi, Shozo (Japan) Kosinski, Antoni A. (USA) Koskela, Pekka J. (Finland) Kosmodemyanskii, Alexander A. (Russian

Federation) Kostant, Ann (USA) Kosugi, Masako (Japan) Kosyak, Alexandr V. (Ukraine) Kota, Osamu (Japan) Kotlyarov, Volodymyr (Ukraine) Kotschick, Dieter (Germany)

Kottwitz, Robert Edward (USA) Koua, Konin (Ivory Coast) Koumandos, Stamatis (Cyprus) Kovács, Sándor J. (USA) Kovalev, Alexander M. (Ukraine) Kovalev, Alexei G. (United Kingdom) Kowalski, Michael (Germany) Kowalski, Piotr (Poland) Kozlov, Vladimir (Sweden) Kracht, Marcus A. (Germany) Krainer, Thomas (Germany) Kraljevic, Hrvoje (Croatia) Krämer, Helmut L. (Germany) Kramer, Jürg (Germany) Kranz, Matthias (Germany) Krasilnikov, Alexei N. (Russian

Federation) Krasnov, Yakov (Israel) Krason, Piotr (Poland) Krattenthaler, Christian (Austria) Kratz, Werner (Germany) Kraus, Anne (Germany) Krause, Henning (Germany) Krause, Matthias (Germany) Krause, Ulrich (Germany) Krebs, Hans-Joachim (Germany) Krejci, Pavel (Germany) Kreußler, Bernd (Germany) Kreuzer, Martin (Germany) Kriecherbauer, Thomas C. (Germany) Krieg, Aloys (Germany) Krikorian, Raphael R. (France) Kriz, Igor (USA) Krolak, Andrzej (Poland) Krötz, Bernhard (Germany) Kruger, Alexander Ya. (Belarus) Kruglikov, Boris Serafimovich (Russian

Federation) Krumke, Sven O. (Germany) Kruse, Hans-Peter (Germany) Krutitskii, Pavel A. (Russian Federation) Kryakin, Yuri (Ukraine) Krzyz, Jan G. (Poland) Kuberek, Monika (Germany) Kubo, Akisato (Japan) Küchler, Uwe (Germany) Kugelmann, Bernd (Germany) Kühn, Ulf (Germany) Kühnel, Arnold (Germany) Kühnel, Wolfgang (Germany)

Kühnlein, Stefan (Germany) Kuksin, Sergei B. (Russian Federation) Kuku, Aderemi O. (Italy) Kulikov, Gennady Yu. (Russian

Federation) Kulikov, Viktor S. (Russian Federation) Kumbhat, R. K. (India) Kummer, Bernd (Germany) Kummetz, Ralph (Germany) Kündgen, André (USA) Kunik, Matthias (Germany) Kunz, Ernst (Germany) Kuo, Tsang-Hai (Taiwan) Kuperberg, Krystyna M. (USA) Kupisch, Herbert (Germany) Kurdachenko, Leonid A. (Ukraine) Kuribayashi, Yukio (Japan) Kurjo McNeill, Sabine U. (United

Kingdom) Kurke, Herbert (Germany) Kürsten, Klaus-Detlef (Germany) Kusch, Claudia (Germany) Küssner, Thilo (France) Kutev, Nicolay Draganov (Germany) Kutz, Nadja (Germany) Kwiecinski, Michal (Poland)

Labarca, Rafael (Chile) Labourie, François (France) Lacey, Michael T. (USA) Lackschewitz, Christina B. (Germany) Laffey, Thomas J. (Ireland) Lafforgue, Laurent (France) Lagarias, Jeffrey C. (USA) Lahtinen, Aatos O. (Finland) Lalonde, Francois (Canada) Lamm, Christoph (Germany) Lamprecht, Erich (Germany) Lance, E. Christopher (United Kingdom) Lanchon, Pierre (France) Lang, Georg (Germany) Lang, Jens (Germany) Lang, Rainer (Germany) Lange, Alexander (Germany) Lange, Otfried (Germany) Lange, Thomas (Germany) Langenbruch, Michael (Germany) Langer, Ottmar (Germany) Langevin, Rémi N. (France) Langmach, Hartmut (Germany) Langmann, Hanns-Heinrich (Germany)

Lanza de Cristoforis, Massimo (Italy) Lanzinger, Hartmut (Germany) Lapidus, Michel L. (USA) Lappan, Peter A. (USA) Lara-Aparicio, Miguel (Mexico) Laradji, Abdallah (Kingdom of

Saudi-Arabia) Lárusson, Finnur (Canada) Lascoux, Alain (France) Lastaria, Federico G. (Italy) Laurenz, Henriette-Marie (Romania) Laurincikas, Antanas (Lithuania) Läuter, Henning (Germany) Lauterbach, Reiner (Germany) Lawrenz, Alexander (USA) Lawton, Wayne M. (Singapore) Lazinska, Anna (Poland) Le, Thang T. Q. (USA) Lebowitz, Joel (USA) Lechner, Patrick O. (Germany) Leck, Uwe (Germany) Leck, Volker (Germany) Leder, Judith (Germany) Ledermann, Walter (United Kingdom) Ledrappier, Francois (France) Lee, Sa Ge (Korea) Lee, Yuh-Jia (Taiwan) Lefton, Lew E. (USA) Le Gall, Jean-François (France) Leger, George F. (USA) Lehalle, Charles-Albert (France) Lehmann, Detlef (Germany) Lehmann, Ingmar (Germany) Lehn, Jürgen (Germany) Lehto, Olli (Finland) Leichtweiß, Kurt (Germany) Leiderman, Arkady (Israel) Leindler, Laszlo (Hungary) Leinert, Michael (Germany) Leitao, Antonio C. (Brazil) Leiterer, Günter Paul (Germany) Leiterer, Jürgen (Germany) Lemaire, Jean-Michel (France) Lemaire-Body, Francoise (France) Lemarié-Rieusset, Pierre Gilles (France) Lempert, Laszlo (USA) Lenard, Christopher T. (Australia) Leng, W. S. (United Kingdom) Lenski, Wolfgang (Germany) Lenz, Daniel H. (Germany)

Lenz, Hanfried (Germany) Lenzing, Helmut (Germany) Leonenko, Nikolai (Ukraine) Leprevost, Franck (France) Leruste, Christian P. (France) Lesch, Matthias (Germany) Leschke, Katrin (Germany) Lescot, Paul E. (France) Leseberg, Dieter (Germany) Lessard, Sabin (Canada) Leti, Guiseppe (Italy) Letzner, Eberhard (Germany) Leutbecher, Armin (Germany) Leviatan, Dany (Israel) Leviatan, Talma O. (Israel) Levin, Alexander (USA) Levin, Norman P. (USA) Levy, Jason (Canada) Levy, Stuart A. (USA) Lewis, Donald J. (USA) Lewis, Wayne (USA) Li, An-Min (P. R. China) Li, Cai-Heng (Australia) Li, Congming (USA) Li, Da-Qian (Ta-Tsien) (P. R. China) Li, Gui-Song (P. R. China) Li, Jiayu (P. R. China) Li, Wei-Ping (P. R. China) Li, Wenlin (P. R. China) Li, Xing (P. R. China) Li, Zhongkai (P. R. China) Liandrat, Jacques (France) Liang, Hua (P. R. China) Liberati, Jose I. (USA) Liberzon, Mark R. (Russian Federation) Lica, Dionis (Romania) Lichtenberg, Heiner (Germany) Lickorish, W. B. Raymond (United

Kingdom) Liebscher, Stefan (Germany) Liemant, Alfred (Germany) Liesen, Jörg (Germany) Lih, Ko-Wei (Taiwan) Lim, Richard (United Kingdom) Lima, Suely dos Santos (Brazil) Lima de Sa, Eduardo (Venezuela) Lin, Zongzhu (USA) Lindblad, Hans (USA) Lindemann, Ina (USA) Lindenstrauss, Joram (Israel)

Lindner, Thomas (Germany)

Ling, San (Singapore) Linton, Fred E. J. (USA) Linton, Stephen A. (United Kingdom) Lipkovski, Aleksandar T. (Yugoslavia) Lipponen, Marjo R. (Finland) Liss, Shari (USA) List, Christoph (Ireland) List, Tobias (Germany) Litvinchuk, Gueorgui S. (Portugal) Litwin, Oleq (Ukraine) Liu, Feng (P. R. China) Liu, Fon-Che (Taiwan) Liu, Zheng (P. R. China) Lizama, Carlos (Chile) Lledó, Fernando (Germany) Löbel, Andreas (Germany) Loday, Jean-Louis (France) Lodder, Gerald M. (USA) Loeffelmann, Jochen (Germany) Loewe, Benedikt (USA) Loewy, Raphael (Israel) Logachev, Dmitry (Russian Federation) Loginov, Boris V. (Russian Federation) Logofet, Dmitrii O. (Russian Federation) Lohkamp, Joachim (Germany) Lommatzsch, Klaus (Germany) Lopes-Pinto, Antonio Jorge Brandao

(Portugal) Lorbeer, Boris (Germany) Lorentz, Rudolph A. (Germany) Loskutov, Alexander (Russian Federation) Losonczi, Attila (Hungary) Lott, John (USA) Lotto, Benjamin A. (USA) Louhivaara, Ilppo Simo (Germany) Love, Eric R. (Australia) Lu, Guozhen (USA) Lu, Sylvia Chin-Pi (USA) Luca, Florian G. (Germany) Luchko, Yuri (Germany) Ludwig, Garry (Canada) Luedecke, Marcus (Germany) Lügger, Joachim (Germany) Luma, Charity (Philippines) Lundmark, Hans (Sweden) Lusala, Tsasa (Germany) Lütkebohmert, Werner (Germany) Lutsenka, Michael (Russian Federation) Lutz, Frank H. (Germany)

Luukkainen, Jouni (Finland) Lyashenko, Elena I. (Ukraine) Lyashenko, Igor N. (Ukraine) Lyons, Jeremiah J. (USA) Lyons, Terry J. (United Kingdom) Lyubeznik, Gennady (USA)

Ma, Li (P. R. China) Ma, Xiaonan (France) Maalouf, Ramez N. (Lebanon) Maass, Peter (Germany) Macdivitt, Arthur (United Kingdom) Macdonald, I. G. (United Kingdom) Macedonska, Olga (Poland) Machedon, Matei (USA) Macintyre, Angus J. (United Kingdom) MacPherson, Robert D. (USA) Maddocks, John H. (Switzerland) Madsen, Dag (Norway) Maeda, Hironobu (Japan) Maehara, Kazuhisa (Japan) Mänz, Oleg (Germany) Magden, Abdullah (Turkey) Magenes, Enrico (Italy) Magid, Andy R. (USA) Magnus, Robert (Iceland) Mahassen, Nadim (United Kingdom) Mahdavi, Kazem (USA) Mahdavi-Hezavehi, Mohammad (Iran) Mahmoodian, Ebadollah S. (Iran) Mahnke, Heidi (Germany) Mahowald, Mark (USA) Maier, Volker (Germany) Mailybaev, Alexei A. (Russian Federation) Makagon, Andrzej (USA) Makarenko, Natalia Yur’evna (Russian

Federation) Makay, Géza (Hungary) Mäkeläinen, Tuulikki (Finland) Makhnev, Alexandre A. (Russian

Federation) Makinde, Oluwole Daniel (Zimbabwe) Malamud, Mark M. (Ukraine) Malik, Grit (Germany) Malik, Surender K. (India) Malinin, Dmitri (Belarus) Mallat, Stéphane (France) Malle, Gunter (Germany) Malliavin, Paul (France) Mallol Gurgui, Josep (Spain) Maltsiniotis, Georges (France)

Mamadaliev, Nazirdjan Kamildjanovich

(Uzbekistan) Mammitzsch, Volker (Germany) Mampassi, Benjamin (Senegal) Manders, Kenneth (USA) Manin, Yuri (Germany) Maniscalco, Caterina (Italy) Manturov, Vasily Olegovich (Russian

Federation) Manuilov, Vladimir (Russian Federation) Marathe, Kishore B. (USA) Marchisio, Marina R. (Italy) Marcolli, Matilde (USA) Marcus, Andrei (Romania) Mardesic, Sibe (Croatia) Marichev, Oleg (USA) Marin, Ivan F. (France) Marinoschi, Gabriela G. (Romania) Marioli, Carlo (Italy) Markarian, Roberto (Uruguay) Markgraf, Norman (Germany) Markina, Irina (Russian Federation) Maroscia, Paolo (Italy) Marrero, Osvaldo (USA) Marshall, Donald E. (USA) Marsico, Tiziana (Italy) Martchenko, Boris G. (Ukraine) Martin, Alexander (Germany) Martin, Benjamin M. (Australia) Martin, Bernd (Germany) Martin-Deschamps, Mireille R. (France) Martinez, Raquel R. (Spain) Martio, Olli T. (Finland) Maruyama, Fumitsuna (Japan) März, Roswitha (Germany) Masanja, Verdiana Grace (Tanzania) Masaya, Mochizuki (Japan) Mascari, Giovanni Francesco (Italy) Maseberg, Sönke (Germany) Maslyuchenko, Volodymyr K. (Ukraine) Massart, Daniel (Germany) Massold, Heinrich (Germany) Materov, Evgeni Nikolaevich (Russian

Federation) Matet, Pierre (France) Mathé, Peter (Germany) Mathieu, Martin (Ireland) Mathieu, Oliver (France) Mathieu, Pierre (Canada) Matoušek, Jiri (Czech Republic)

Matsumoto, Shigeki (Japan) Matsumoto, Shigenori (Japan) Matsumoto, Waichiro (Japan) Matsumura, Toru (Japan) Matsuyama, Yoshio (Japan) Matsuzaki, Nagi (Japan) Matthies, Karsten (Germany) Mattila, Pertti (Finland) Mattner, Lutz (Germany) Mattrisch, Sybille (Germany) Matuszewski, Roman (Poland) Matveev, Oleg V. (Russian Federation) Matveev, Vladimir S. (Russian

Federation) Maude, Ronald (United Kingdom) Maumary, Serge (Switzerland) Maxwell, James W. (USA) Mayer, John C. (USA) Mayer, Karl Heinz (Germany) Mbakop, Guy Merlin (Germany) Mbunga, Paulo (Germany) McCoy, Barry (USA) McCune, Catherine (Germany) McDuff, Dusa (USA) McMullen, Curtis Tracy (USA) Mc Nulty, Jennifer (USA) Mdzinarishvili, Leonard (Georgia) Medghalchi, Alireza (Iran) Mednykh, Alexander D. (Russian

Federation) Megginson, Robert E. (USA) Mehri, Bahman (Iran) Mehta, Ghanshyam B. (Australia) Meier, David (Switzerland) Meintrup, David (Germany) Meissner, Nils (Germany) Meltzer, Hagen (Germany) Melville, Duncan J. (USA) Mendes, Diana E. Aldea (Portugal) Mendez, Jose (Germany) Mendoza, Ramon (Brazil) Mennicken, Reinhard (Germany) Mercier, Armel (Canada) Merel, Loic (France) Merklen, Héctor (Brazil) Merle, Frank Eric (France) Mertins, Ulrich (Germany) Merzon, Anatoli (Russian Federation) Meskhi, Alexander N. (Georgia) Mestrovic, Romeo (Yugoslavia)

Metsänkylä, Tauno (Finland) Metzger, Roger J. (Brazil) Meusers, Claudia (Germany) Meyer, Gottfried Peter (Germany) Meyer, Johannes H. (South Africa) Meyer, Karoline (Germany) Miana, Pedro J. (Spain) Miatello, Roberto J. (Argentina) Michaelis, Walter J. (USA) Michailov, Leonid Grigorevich (Tajikistan) Micheel, Jan (Germany) Michel, Alexis P. (France) Michel, Volker (Germany) Michelucci, Maria Letizia (Italy) Michev, Iordan P. (Japan) Michor, Peter W. (Austria) Micula, George (Romania) Mielke, Hans F. (Germany) Migda, Malgorzata (Poland) Mignanego, Fausto (Italy) Migorski, Stanislaw (Poland) Mihalache, George (Germany) Mihalache, Nicolae S. (Romania) Mihalyko, Csaba (Hungary) Mikaelian, Vahagn H. (Armenia) Mikhalev, Alexander A. (Russian

Federation) Mikhalev, Alexandre V. (Russian

Federation) Mikhalkin, Grigory (USA) Milka, Anatoly D. (Ukraine) Miller, Stephen D. (USA) Millett, Kenneth C. (USA) Millionshikov, Dmitri V. (Russian

Federation) Milman, Vitali (Israel) Milnes, Paul (Canada) Milnor, John W. (USA) Milstein, Grigori (Germany) Milton, Graeme Walter (USA) Minda, C. David (USA) Minian, Elias G. (Germany) Mio, Washington (USA) Mirenghi, Elvira (Italy) Misaki, Norihiro (Japan) Mishchenko, Alexander S. (Russian

Federation) Mishura, Yuliya Stepanovna (Ukraine) Misiewicz, Jolanta K. (Poland) Misiolek, Gerard (USA)

Misra, Kailash C. (USA) Missarov, Moukadas (Russian Federation) Mititelu, Stefan M. (Romania) Mitran, Ilie (Romania) Mitschi, Claude (France) Mitsumatsu, Yoshihiko (Japan) Mittenhuber, Dirk (Germany) Miwa, Tetsuji (Japan) Miyajima, Kimio (Japan) Mizutani, Tadayoshi (Japan) Mlynarska, Elzbieta (Poland) Mochizuki, Shinichi (Japan) Moffatt, Keith (United Kingdom) Mogk, Georg (Germany) Mohammadzadeh, Mohsen (Iran) Mohammed, Alip (Germany) Mohn, Karl-Heinz (Germany) Mohnke, Klaus (Germany) Möhring, Rolf H. (Germany) Moklyachuk, Mikhail (Ukraine) Moldavskaya, Elina M. (Ukraine) Möller, Andris (Germany) Möller, Herbert (Germany) Monastyrsky, Michael (Russian

Federation) Monod, Nicolas (Switzerland) Montans, Fernando J. (Uruguay) Montgomery, M. Susan (USA) Moore, Charles N. (Ireland) Moore, John D. (USA) Morales Figueroa, Bernardo (Guatemala) Morawetz, Cathleen Synge (USA) Moree, Pieter (Germany) Moreira, Carlos C. (Brazil) Moreira, Carlos Gustavo T. de A. (Brazil) Morgenthal, Klaus (Germany) Morgulis, Andrey B. (Russian Federation) Mori, Masao (Japan) Morimoto, Hiroko (Japan) Morimoto, Mitsuo (Japan) Moritoh, Shinya (Japan) Morley, Mary E. (USA) Moroianu, Andrei, M. (France) Morrell, Robert (United Kingdom) Morris, Alun O. (United Kingdom) Mörters, Peter (Germany) Moser, Jürgen (Switzerland) Moshnikova, Julia M. (Russian

Federation) Motohashi, Tomoe (Japan)

Motornaya, Oxana V. (Ukraine) Mourrain, Bernard (France) Moussa, Gonoko (France) Moussavi, Ahmad (Iran) Mouton, Sonja (South Africa) Movisyan, Moevses (Armenia) Mozes, Shahar (Israel) Mucha, Hans-Joachim (Germany) Mueller, Peter (Germany) Mueller-Roehlck, Karin (Germany) Müger, Michael (Italy) Mulholland, Lawrence S. (United

Kingdom) Müller, Detlef H. (Germany) Müller, Eva Nuria (Germany) Müller, Franz (Switzerland) Müller, Stefan (Germany) Müller-Hannemann, Matthias (Germany) Mumford, David (USA) Mündlein, Heidrun (Germany) Munembe, Joao S. P. (Mozambique) Munkholm, Hans J. (Denmark) Murakami, Hitoshi (Japan) Murali, Venkat (South Africa) Muranov, Yuri V. (Russian Federation) Murasugi, Kunio (Canada) Muscalu, Florin C. (USA) Mushtari, Daniar H. (Russian Federation) Mustonen, Vesa (Finland) Myshkis, Anatoli (Russian Federation)

Näätänen, Marjatta (Finland) Nagaev, Sergei (Russian Federation) Nagahara, Takashi (Japan) Nagaoka, Shoyu (Japan) Nagasaka, Kenji (Japan) Nagasawa, Takeyuki (Japan) Nagel, Uwe (Germany) Nagy, Gábor P. (Hungary) Näher, Holger D. (Germany) Nair, Sunil (United Kingdom) Naito, Hirotada (Japan) Nakagawa, Jin (Japan) Nakai, Eiichi (Japan) Nakamura, Iku (Japan) Nakamura, Kirio K. (Japan) Nakano, Daniel K. (USA) Nakano, Minoru (Japan) Nakayashiki, Atsushi, An (France) Nakazawa, Hideaki (Japan) Nanbu, Tokumori (Japan)

Narasimhan, Mudumbai Seshachalu

(Italy) Narayanaswami, Pallasena P. (Canada) Nassau, Christian (Germany) Nassopoulos, Georgios F. (Greece) Nastold, Hans-Joachim (Germany) Nasyrov, Samyon R. (Russian Federation) Natanzon, Sergei (Russian Federation) Natarajan, Raja (India) Natroshvili, David (Georgia) Natsume, Toshikazu (Japan) Naude, Cornelia G. (South Africa) Naumann, Joachim (Germany) Naveira, Antonio M. (Spain) Naydonova, Anna V. (Ukraine) Nazarova, Ludmila Aleksandrovna

(Ukraine) Nazipov, Albert A. (Russian Federation) Ndimubandi, Jean (Burundi) Neeb, Karl-Hermann (Germany) Nefedov, Nikolay N. (Russian Federation) Neff, Patrizio (Germany) Nejaddeghan, Yadollah (Iran) Nelius, Christian-Frieder (Germany) Nepotu, Gabriel-Lucian (Romania) Nesetril, Jaroslav (Czech Republic) Nesselmann, Dieter (Germany) Neto, Orlando (Portugal) Netsvetaev, Nikita Yu. (Russian

Federation) Neubauer, Mathias (Germany) Neuberger, Barbara O. (USA) Neuberger, John W. (USA) Neumann, Bernhard H. (Australia) Neumann, Heike (Germany) Neumann, Olaf G. (Germany) Neumann, Peter (Germany) Neun, Winfried (Germany) Neunhaeuserer, Joerg (Germany) Neusel, Mara D. (Germany) Newelski, Ludomir (Poland) Nezit, Pierre (Ivory Coast) Ngom, Papa (France) Nguyen, Anh-Tuan (Germany) Nguyen, Dinh Cong (Vietnam) Nguyen, Dinh Tri (Vietnam) Ngwa, Gideon A. (Cameroon) Nicodeme, Pierre (Germany) Nicolae, Florin (Romania) Nicolas, Anne-Marie (France)

Nicolas, Jean-Louis (France) Niebsch, Jenny (Germany) Niederhauser, Beat M. (Germany) Niederkorn, Philippe (Belgium) Niederreiter, Harald (Austria) Nierhoff, Till (Germany) Nieto, Isidro P. (Mexico) Nietzsch, Josef (Germany) Niezgoda, Marek (Poland) Nikcevic, Stana (Yugoslavia) Nikitin, Igor N. (Germany) Nikolova, Ivanka P. (USA) Nikoltjeva-Hedberg, Margarita (Sweden) Nikonov, Oleg I. (Russian Federation) Nill, Benjamin (Germany) Nilsson, Marcus L. G. (Sweden) Ninnemann, Olaf (Germany) Ninomiya, Yasushi (Japan) Nishimura, Jun-ichi (Japan) Nishimura, Shigeto (Japan) Niss, Mogens (Denmark) Noaghi, Sorin (Romania) Nocedal, Jorge (USA) Nolan, Clifford J. (USA) Nollau, Volker (Germany) Nomura, Yasutoshi (Japan) Nongxa, Loyiso (South Africa) Nörenberg, Rainer (Germany) Norkin, Vladimir (Ukraine) Nörtemann, Stefan (Germany) Nourani, Cyrus F. (USA) Novikov, Igor Ya. (Russian Federation) Nowack, Dieter (Germany) Nowak, Ivo (Germany) Nowak, Maria T. (Poland) Nowak, Ulrich (Germany) Nummelin, Esa (Finland) Nürnberg, Reiner (Germany) Nüsken, Michael (Germany) Nussbaum, Michael (Germany) Nusser, Harald (Germany)

Oakes, Susan M. (United Kingdom) Oberschelp, Arnold (Germany) Obukhovskii, Valeri V. (Russian

Federation) Odai, Yoshitaka (Japan) Odell, Edward (USA) Oevel, Walter (Germany) Oeverdieck, Lars (Germany) Ogata, Hayao (Japan)

Ogata, Shoetsu (Japan) Ohba, Kiyoshi (Japan) Ohmiya, Mayumi,EM. (Japan) Ohno, Hiloshi (Japan) Ohno, Masahiro (Japan) Ohno, Shuichi (Japan) Ohta, Minolu (Japan) Ohtsuki, Tomotada (Japan) Oja, Eve (Estonia) Oja, Peeter (Estonia) Oka, Hiro-e (Japan) Okada, Tatsuya (Japan) Okamoto, Hisashi (Japan) Okayasu, Takashi (Japan) Okikiolu, Kate (USA) Okoya, Samuel S. (Nigeria) Olbrich, Martin (Germany) Oleinik, Vladimir L. (Russian Federation) Olenko, Andrew Ya. (Ukraine) Oliver, Robert A. (France) Olshanskii, Maxim A. (Russian

Federation) Olsson, Jorn B. (Denmark) Omarjee, Moubinool (France) O’Neill, Bruce (Egypt) Onishi, Yoshihiro (Japan) Ono, Kaoru (Japan) Ono, Ken (USA) Ontaneda, Pedro (Brazil) Opfer, Gerhard (Germany) Opozda, Barbara (Poland) Orazov, Mered (Turkmenistan) Orman, Gabriel V. (Romania) O’Shea, Donal B. (USA) Osilike, Micah O. (Nigeria) Osipenko, George (Russian Federation) Ostermann, Alexander (Austria) Osthus, Deryk (Germany) Otachel, Zdzislaw (Poland) Otsuka, Kayo (Japan) Otsuka, Kenichi (Japan) Ovchinnikov, Vladimir I. (Russian

Federation) Oversteegen, Lex G. (USA) Owa, Shigeyoshi (Japan) Özlük, Ali E. (USA) Paatz, Andrea (Germany) Paccagnan, Diego (Italy) Pacheco Esteban, Juan Pablo (Spain) Padial, Juan Francisco (Spain)

Paditz, Ludwig (Germany) Paetsch, Olaf (Germany) Paffenroth, Randy (Switzerland) Pahlig, Hans-Christian (Germany) Palais, Richard, S. (Germany) Palamodov, Victor P. (Germany) Palencia, Cesar (Spain) Palis, Jacob (Brazil) Palka, Bruce P. (USA) Panazzolo, Daniel C. (Brazil) Pandey, Jagdish N. (Canada) Panov, Evgeniy Yur’evich (Russian

Federation) Pantelidis, Georgios (Greece) Papanicolaou, George C. (USA) Papaschinopoulos, Garyfalos (Greece) Pappas, Christian E. (USA) Parkhomenko, Victoria (Ukraine) Parmenter, Michael M. (Canada) Parring, Aivo (Estonia) Parthasaraty, Kalyanapuram Rangachari

(India) Paseman, Gerhard R. (USA) Passaquindici, Maria (Italy) Passare, Mikael (Sweden) Patetta, Nicolás D. (Argentina) Pauen, Raimund (Germany) Pauly, Marc (Belgium) Pavcevic, Mario-Osvin (Croatia) Pavel, Monique L. (France) Pavicevic, Zarko (Yugoslavia) Pawalowski, Krzysztof M. (Poland) Pawlowsky-Glahn, Vera (Spain) Paxia, Giuseppe (Italy) Paycha, Sylvie (France) Pecher, Hartmut (Germany) Pedas, Arvet (Estonia) Pedit, Franz (Germany) Pein, Jan (Germany) Pekonen, Osmo (Finland) Pelczar, Andrzej (Poland) Peller, Vladimir V. (USA) Peltonen, Kirsi (Finland) Penssel, Christian (Germany) Perepelitsa, Vitaly A. (Russian

Federation) Peretyat’kin, Mikhail (Kazakhstan) Pereverziev, Sergii (Ukraine) Perez, Javier (Spain) Perez Riera, Mario (Spain)

Peric, Veselin S. (Yugoslavia) Perisic, Vesna (Germany) Perov, Anatolii Ivanovich (Russian

Federation) Perovic, Miodrag (Yugoslavia) Persson, Lars-Erik (Sweden) Pesch, Hans Josef (Germany) Pete, Gabor (Hungary) Petean, Jimmy (Germany) Peters, Klaus (USA) Peters, Meinhard H. (Germany) Petersen, Carsten Lunde (Denmark) Petersson, Henrik (Sweden) Petrogradsky, Victor M. (Russian

Federation) Petrosjan, Leon A. (Russian Federation) Petti, Richard J. (USA) Petukhov, Alexander P. (Russian

Federation) Petzoldt, Martin (Germany) Peyre, Emmanuel (France) Pfeiffer, Helmut (Germany) Pfeiffer, Ruth (USA) Pfender, Michael (Germany) Pfetsch, Marc E. (Germany) Pfister, Albrecht (Germany) Pfister, Gerhard (Germany) Pflaum, Christoph (Germany) Pflaum, Markus J. (Germany) Pflueger, Klaus (Germany) Philip, Peter (Germany) Philippin, Gerard A. (Canada) Phillips, Keith L. (USA) Phillips, N. Christopher (USA) Phung, Ho (Vietnam) Pickrell, Douglas M. (USA) Piefke, Frank (Germany) Piene, Ragni (Norway) Pier, Jean-Paul (Luxembourg) Pilipovic, Stevan (Yugoslavia) Pillay, Anand (USA) Pilyugin, Sergei Yu. (Russian Federation) Pilz, Alexander (Germany) Pinchuk, Sergey Ivanovich (Russian

Federation) Pinheiro, Vilton (Brazil) Pink, Richard (Germany) Pinkall, Ulrich (Germany) Pinto, Alberto Adrego (Portugal) Piovan, Luis A. (Argentina)

Pisier, Gilles (France) Pitassi, Toniann (USA) Pittet, Christophe (Germany) Pizzo, Aldo Bruno (Argentina) Plato, Robert (Germany) Platonov, Sergei S. (Russian Federation) Platzek, Maria I. (Argentina) Plaza, Sergio (Chile) Plewe, Klaus J. (Germany) Plümer, Judith (Germany) Poensgen, Diana A. I. (Germany) Pöhle, Bärbel (Germany) Pöhle, Uwe (Germany) Pohozaev, Stanislav I. (Russian

Federation) Pohst, Michael E. (Germany) Poldvere, Märt (Estonia) Polterovich, Leonid (Israel) Polthier, Konrad (Germany) Polyakova, Katerina (Russian Federation) Polzehl, Jörg (Germany) Pomazanov, Michael V. (Russian

Federation) Ponce, Gustavo Alberto (USA) Ponge, Raphael S. (France) Ponomarenko, Andrej (Germany) Ponomarev, Paul (USA) Popa, Anca (Romania) Popescu, Dorin-Mihail (Romania) Popov, Igor Yu. (Russian Federation) Popova, Elena V. (Russian Federation) Porteous, Hugh L. (United Kingdom) Poschadel, Norbert (Germany) Pöschel, Reinhard (Germany) Post, Katharina D. E. (Germany) Postelnicu, Tiberiu V. (Romania) Postnikov, Mihkail M. (Russian

Federation) Potapov, Vadim D. (Russian Federation) Potthast, John (USA) Potthast, Roland (Germany) Pourkazemi, Mohammad H. (Iran) Pragarauskas, Henrikas (Lithuania) Präve, Alexandra (Germany) Preissler, Gabi (Germany) Presman, Ernst L. (Russian Federation) Prestin, Jürgen (Germany) Preston, Gordon B. (Australia) Preunkert, Marc (Germany) Preuß, Gerhard (Germany)

Pribitkin, Wladimir de Azevedo

(Germany) Priebe, Volker (Germany) Prieto, Carlos (Mexico) Priplata, Christine (Germany) Prishlyak, Alexander (Ukraine) Prokhorov, Yuri (Russian Federation) Prömel, Hans J. (Germany) Prüfer, Friedbert (Germany) Pruschke, Thilo (Germany) Prytula, Yaroslav Ya. (Ukraine) Ptak, Marek (Poland) Puchta, Jan-Christoph (Hungary) Pujals, Enrique Ramiro E. R. P. (Brazil) Pukhlikov, Aleksandr (Russian

Federation) Pulkina, Ludmila S. (Russian Federation) Pulleyblank, William R. (USA) Puppe, Dieter (Germany) Puppe, Volker (Germany) Pustylnikov, Lev L. D. (Russian

Federation)

Quast, Peter (Germany) Quian, Minping (P. R. China) Quine, John R. (USA) Quinn, Frank (USA)

Rabindranathan, Malayattil (USA) Rabinovich, Vladimir S. (Russian

Federation) Rachele, Lizabeth V. (USA) Rack, Heinz-Joachim (Germany) Racke, Reinhard (Germany) Rackwitz, Hans-Georg (Germany) Raczynski, Andrzej (Poland) Rademacher, Hans-Bert (Germany) Rademacher, Jens (Germany) Radic, Mirko V. (Sweden) Radtke, Achim (Germany) Radu, Mircea (Romania) Radyno, Alexander Yakovlevich (Belarus) Radyno, Nikolai Ya. (Belarus) Raghavan, K. N. (India) Raghavendra, Nyshadham (India) Ragusa, Alfio (Italy) Raikov, George D. (Bulgaria) Raitums, Uldis (Latvia) Rajabov, Nusrat (Tajikistan) Rakhimov, Abdugafaz (Uzbekistan)

Rakic, Zoran (Yugoslavia) Rakosník, Jirí (Czech Republic) Ramachandran, Niranjan (USA) Rambau, Jörg (Germany) Ramirez, Alejandro F. (Switzerland) Ramirez-De-Arellano, Enrique (Mexico) Ramollo, Motlatsi Pius (Lesotho) Randow von, Thomas (Germany) Rangarajan, Govindan (India) Rapoport, Michael (Germany) Rappoport, Juri M. (Russian Federation) Rassias, Themistocles M. (Greece) Rasskazov, Alexey (Russian Federation) Rathsfeld, Andreas (Germany) Ratnakumar, Peetta Kandy (India) Rawat, Rama (Israel) Rayskin, Victoria (USA) Rebelo, Julio Cesar (Brazil) Rebholz, Joachim (United Kingdom) Rebiai, Salah-Eddine (Algeria) Rebolledo, Rolando (Chile) Recke, Lutz (Germany) Reddy, B. Daya (South Africa) Reemtsen, Rembert (Germany) Rees, S. Mary (United Kingdom) Rego, F. Alfredo (Guatemala) Rehberg, Joachim (Germany) Rehmann, Ulf (Germany) Reichel, Hans-Christian (Austria) Reichert-Hahn, Marianne (Germany) Reigle, Dennis J. (USA) Reiher, Gesine (Germany) Reimann, Hans Martin (Switzerland) Reinelt, Gerhard (Germany) Reinhardt, Gerd (Germany) Reiris i Thurralde, Martin (Uruguay) Reiss, Markus (Germany) Reiten, Idun (Norway) Reitz, Stefan (Germany) Rempala, Grzegorz A. (USA) Ren, Nanheng (P. R. China) Rendall, Alan D. (Germany) Renteria-Marquez, Carlos (Mexico) Rentner, Wolf-Dietrich (Germany) Rentschler, Rudolf (France) Repovs, Dusan (Slovenia) Rerikh, Konstantin V. (Russian

Federation) Ressel, Paul (Germany) Revalski, Julian P. (Bulgaria)

Reyes Rodríguez, Armando (Mexico) Riahi, Hasna (Tunesia) Ribe, Martin G. (Sweden) Ricci, Fulvio (Italy) Richter, Thomas (Germany) Rickard, Jeremy C. (United Kingdom) Rickmann, Seppo U. (Finland) Riede, Adolf J. I. (Germany) Rieder, Andreas (Germany) Rieder, Ulrich (Germany) Riedl, Johannes (Germany) Rieger, Marc Oliver (Germany) Riemenschneider, Oswald W. (Germany) Riesel, Hans (Sweden) Rieutord, Michel L. E. (France) Rindler, Harald (Austria) Ringel, Claus Michael (Germany) Riordan, Oliver M. (United Kingdom) Ritter, Günter (Germany) Rizvi, Syed Tariq (USA) Robbiani, Marcello L. (Switzerland) Robbiano, Lorenzo (Italy) Robert, Aline (France) Roberts, Brooks K. (Canada) Robertson, Ian (USA) Robinson, Bencion (Israel) Robinson, Margaret M. (USA) Rocsoreanu, Carmen (Romania) Roczen, Marko (Germany) Rodkina, Alexandra E. (Russian

Federation) Rodrigues, Jose Francisco (Portugal) Rodriguez del Rio, Roberto (Spain) Rodriguez Santiesteban, Antonio Ramon

(Germany) Rodriguez Trueba, Maria Isabel (Spain) Roegner, Katherine S. (Germany) Roendigs, Oliver (Germany) Rögen, Peter (Denmark) Rogmann, Sascha A. (Germany) Rogowski, Andrzej (Poland) Roitzsch, Rainer (Germany) Rolles, Silke W. (Germany) Roman, Olena (Ukraine) Römisch, Werner (Germany) Ronconi, Maria Cristina (Italy) Ronto, Miklos (Hungary) Roos, Jan-Erik I. (Sweden) Roque, Tatiana (France) Rosa, Renata (Brazil)

Rosado, Fernando (Portugal) Rosemeier, Frank J. (Germany) Rosenbaum, Kurt (Germany) Rosenbaum, Wolfram (Germany) Rosenberger, Gerhard (Germany) Rosenknop, John (Germany) Rosenthal, Haskell P. (USA) Rösler, Margit (Germany) Rosset, Shmuel (Israel) Rossetti, Juan Pablo (Argentina) Rossi, Louis F. (USA) Rothschild, Linda P. (USA) Roubicek, Tomas (Czech Republic) Rouillier, Fabrice F. R. (France) Roushon, Sayed Khaled Mr. (Germany) Rowe, David E. (Germany) Rowley, Christopher (United Kingdom) Royter, Andriy (Ukraine) Rozanova, Olga S. (Russian Federation) Rozanski, Roman (Poland) Ru, Min (USA) Ruan, Yongbin (USA) Rubin, Boris (Israel) Ruckmann, Jan-J. (Germany) Rudolf, Ingeborg (Germany) Rudyak, Yuli B. (Germany) Ruffing, Andreas (Germany) Ruh, Ernst A. (Switzerland) Runge, Bernhard (Germany) Runge, Daniel (Germany) Rusch, Beate (Germany) Rüssmann, Helmut (Germany) Ruzicka, Michael (Germany) Ryska, Norbert (Germany) Ryushi, Goto (Japan)

Sabitova, Mary Nailevna (Russian

Federation) Sablatnig, Jan-H. (Germany) Sadyrbaev, Felix (Latvia) Safonov, Mikhail Vasilevich (USA) Sahab, Salem A. (Kingdom of

Saudi-Arabia) Saintloup, Bernard (France) Saitoh, Hitoshi (Japan) Sakakibara, Nobuhisa (Japan) Sakata, Hiroshi (Japan) Saktikov, Evgueni (Russian Federation) Salberger, Per (Sweden) Saldanha, Nicolau (Brazil) Salinas, Norberto (USA)

Salinger, David L. (United Kingdom) Salmhofer, Manfred (Switzerland) Sambarino Ottino, Martin Javier

(Uruguay) Samoilenko, Anatolij M. (Ukraine) Sanatani, Sunoy (Canada) Sanchez Ruiz, Luis M. (Spain) Sancho, Fernando (Spain) Sand, Guido (Germany) Sandler, Hanna M. (USA) Saneblidze, Samson (Georgia) Sangare, Daouda (France) San Martín, Bernardo A. (Chile) Sanni, Sikiru A. (Nigeria) Sano, Kimiro (Japan) Santos, Francisco (Spain) Sanz-Solé, Marta (Spain) Sarhan, Ammar (Egypt) Sarnak, Peter (USA) Sasao, Seiya (Japan) Sastry, Seemanapalli V. (Brazil) Satake, Ichiro (Japan) Sato, Atsushi (Brazil) Sato, Ken-iti (Japan) Sato, Yumiko (Japan) Sauermann, Markus (Germany) Savas, Ekrem (Turkey) Saveliev, Nikolai (USA) Sawada, Ken (Japan) Sawae, Ryuichi (Japan) Sawon, Justin (United Kingdom) Sawyer, Patrice (Canada) Sax, Ulrich K. (Germany) Sbordone, Carlo (Italy) Scardua, Bruno (Brazil) Scepanovic, Radoje (Yugoslavia) Schachermayer, Walter (Austria) Schäl, Manfred (Germany) Scharlach, Christine (Germany) Scharlau, Rudolf M. E. (Germany) Scharlau, Winfried (Germany) Scharlemann, Martin (USA) Schauenburg, Peter (Germany) Schechtman, Gideon (Israel) Scheel, Arnd (Germany) Scheerer, Hans (Germany) Scheiderer, Claus (Germany) Schein, Boris M. (USA) Schendel, Udo (Germany) Schenzel, Peter (Germany)

Schermuly, Carola M. (Germany) Scheutzow, Michael K. R. (Germany) Schied, Alexander (Germany) Schiffels, Gerhard (Germany) Schilling, Susan (USA) Schilling, Tanja (Germany) Schimmerling, Ernest (USA) Schimming, Rainer (Germany) Schinas, Christos J. (Greece) Schinas, John (Greece) Schleicher, Dierk (Germany) Schlesinger, Karl-Georg (Germany) Schlichting, Günter H. (Germany) Schlichting, Marco (France) Schlickewei, Hans Peter (Germany) Schlippe, Johannes (Germany) Schlomiuk, Dana (Canada) Schlosser, Hartmut (Germany) Schlundt, Rainer (Germany) Schmalz, Gerd (Germany) Schmeelk, John (USA) Schmehr, Dirk R. W. (Germany) Schmeling, Joerg (Germany) Schmelzer, Ilja (Germany) Schmickler-Hirzebruch, Ulrike (Germany) Schmid, Wilfried (USA) Schmidt, Alexander (Germany) Schmidt, Andreas U. (Germany) Schmidt, Frank (Germany) Schmidt, Gunther (Germany) Schmidt, Marco (Germany) Schmidt, Rita Maria (Germany) Schmidt, Siegbert (Germany) Schmidt, Stefan E. (USA) Schmidt, Werner (Germany) Schmidt-Ehrenberg, Johannes (Germany) Schmidtmann, Olaf (Germany) Schmieta, Stefan H. (USA) Schmitt, Burkhard J. (Germany) Schmitt, Nikolaus (Germany) Schmitt, Peter (Austria) Schmüdgen, Konrad (Germany) Schneider, Kai (Germany) Schneider, Klaus R. (Germany) Schnürer, Oliver C. (Germany) Schock, Eberhard (Germany) Schoenwaelder, Ulrich F. K. (Germany) Scholz, Hartmut (Germany) Scholz, Reinhard (Germany) Schönhage, Arnold (Germany)

Schonmann, Roberto (USA) Schoppmeier, Ulrich (Germany) Schork, Matthias (Germany) Schrader, Rainer (Germany) Schrader, Robert (Germany) Schramm, Ruben (Israel) Schreyer, Frank (Germany) Schrijver, Alexander (Netherlands) Schrohe, Elmar (Germany) Schubert, Roman (Germany) Schult, Jörg (Germany) Schulte, Udo (Germany) Schultz, Rüdiger (Germany) Schultz, Reinhard E. (USA) Schulz, Andreas S. (Germany) Schulz, Bernd (Germany) Schulz, Christian (Germany) Schulz, Dirk (Germany) Schulz, Ralph-Hardo (Germany) Schulz, Ursula (Germany) Schulz, Wolfgang (Germany) Schulz-Baldes, Hermann (Germany) Schulze, Bert-Wolfgang (Germany) Schulze, Felix (Germany) Schulze, Volker (Germany) Schulze-Pillot, Rainer (Germany) Schurath, Ronald (Germany) Schurr, Ingo (Germany) Schurz, Henri U. (Colombia) Schuster, Peter Michael (Germany) Schwachhöfer, Lorenz (Germany) Schwartz, Alan L. (USA) Schwartz, Alexander (Germany) Schwartz, Thomas (Germany) Schwarz, Ernst (Germany) Schwarz, Wolfgang K. (Germany) Schweigert, Dietmar (Germany) Schweizer, Jürgen (Germany) Schweizer, Martin (Germany) Schwenk, Angela (Germany) Schwertfeger, Karl-Heinz (Germany) Schwichtenberg, Helmut (Germany) Schwitzer, Felix (Austria) Scott, Richard (USA) Sczech, Robert (USA) Sedykh, Vyacheslav D. (Russian

Federation) Seeländer, Jörg (Germany) Seidel, Jacob (Netherlands) Seifoullaev, Roustam K. (Azerbaijan)

Seiler, Joerg (Germany) Seiler, Ruedi (Germany) Seiler, Wolfgang K. (Germany) Seip, Kristian (Norway) Sekiguchi, Takeshi (Japan) Sekita, Eitaro (Japan) Selder, Erich (Germany) Selfridge, John L. (USA) Seligman, George B. (USA) Selivanova, Elena Nikolaevna (Russian

Federation) Selwat, Karol (Poland) Semeria, Severine (France) Semmler, Beate (Germany) Semmler, Klaus-Dieter (Switzerland) Senf, Peter (Germany) Serbetcioglu, Aglaia Marissa (Germany) Serganova, Vera (USA) Sergeev, Armen (Russian Federation) Series, Caroline (United Kingdom) Sevenster, Arjen (Netherlands) Seydel, Detlef (Germany) Sezinando, Helena (Portugal) Shabozov, Mirgand (Tajikistan) Shadman, Darinsh (Iran) Shadwick, William F. (United Kingdom) Shah, Nimish A. (India) Shah, Riddhi (India) Shahvarani-Semnani, Ahmad (Iran) Shaidurov, Vladimir V. (Russian

Federation) Shaikhet, Leonid E. (Ukraine) Shakunle, Lere O. (Germany) Shalev, Aner (Israel) Shaneson, Julius L. (USA) Sharipbaev, Altynbek (Kazakhstan) Sharma, B. K. (India) Shatz, Stephen S. (USA) Shaw, Sen-Yen (P. R. China) Shchepanika, Elena (Russian Federation) Shchepanuk, Gennady (Ukraine) Shen, Yu (USA) Shepelsky, Dmitry (Ukraine) Sheshko, Michal (Poland) Sheth, Dilip (D. N.) (India) Shevchenko, Yurii (Russian Federation) Shevelev, Vladimir S. (Russian

Federation) Shi, Niandong (USA) Shi, Wujie (P. R. China)

Shiba, Hironori (Japan) Shibano, Hiroki (Japan) Shibata, Tetsutaro T. (Japan) Shima, Kazuhisa (Japan) Shimomura, Katsunori (Japan) Shimura, Tatsuya (Japan) Shintani, Toshitada (Japan) Shiraiwa, Kenichi K. S. (Japan) Shiratani, Katsumi (Japan) Shirikyan, Armen (Russian Federation) Shishikura, Mitsuhiro (Japan) Shiu, Christine (United Kingdom) Shiu, Peter (United Kingdom) Shmygewskyi, Mykola (Ukraine) Shoji, Mayumi (Japan) Shomron, Noam (USA) Shor, Naum Zuselevich (Ukraine) Shor, Peter W. (USA) Shtern, Alexander I. (Russian Federation) Shtykel, Tatyana (Germany) Shu, Chi-Wang (USA) Shubin, Mikhail (USA) Shubladze, Mamuka (Georgia) Siddo, Carol (United Kingdom) Siebeneicher, Christian (Germany) Siebenmann, Laurent C. (France) Sieber, Jan (Germany) Siegfanz, Monika (Germany) Siegmund, David Oliver (USA) Siegmund, Stefan (Germany) Siegmund-Schultze, Rainer (Germany) Siegmund-Schultze, Reinhard (Germany) Siemon, Helmut (Germany) Sieveking, Malte (Germany) Sigmund, Karl (Austria) Sigrist, Francois (Switzerland) Sigurdsson, Ragnar (Sweden) Silva, Jorge-Nuno O. (Portugal) Simeliene, Nijole (Lithuania) Simic, Slavko (Yugoslavia) Simic, Slobodan (USA) Simic, Slobodan K. (Yugoslavia) Simms, David J. (Ireland) Simon, Jacques C. H. (France) Simon, Udo H. (Germany) Simons, Gordon E. (Canada) Sims, Charles C. (USA) Singer, Ivan (Romania) Singh, Ajit Iqbal (India) Singh, Anand Prakashi (India)

Singh, Shobh N. (South Africa) Sireteanu, Tudor (Romania) Siu, Man-Keung (P. R. China) Sjögren, Peter (Sweden) Skill, Thomas (Germany) Skopenkov, Arkady (Russian Federation) Skrypnik, Igor I. (Ukraine) Skutella, Martin (Germany) Slateff, Andreas (Austria) Slattery, Suzanne (USA) Slawig, Thomas (Germany) Slembek, Silke (France) Sloane, Neil J. A. (USA) Slovak, Jan (Czech Republic) Slungård, Inger Heidi (Norway) Smale, Steve (P. R. China) Smirnov, Fedor (France) Smirnov, Georgi V. (Portugal) Smirnov, Stanislav (Germany) Smital, Jaroslav (Czech Republic) Smith, David Alexander (USA) Smith, Hart F. (USA) Smith, Perry B. (USA) Smith, Stuart P. (USA) Smoluk, Antoni (Poland) Soares, Marcio G. (Brazil) Sobolev, Vladimir (Russian Federation) Soboleva, Eugenia S. (Russian Federation) Socolowsky, Jürgen (Germany) Soergel, Wolfgang (Germany) Sofronidis, Nikolas Efstathiou (USA) Soibelman, Yan (USA) Sokolowski, Jan (France) Solano, Cesar (Germany) Solberg, Oyvind (Norway) Söllner, Hildegard (Germany) Solodkyy, Sergiy Georgievich (Ukraine) Sologuren, Santiago (Bolivia) Sommerhäuser, Yorck (Germany) Sonin, Constatin I. (Russian Federation) Sonntag, Rolf (Germany) Sormani, Christina (USA) Sos, Vera T. (Hungary) Sostak, Alexander (Latvia) Soto-Andrade, Jorge (Chile) Soulé, Christophe (France) Spandaw, Jeroen G. (Germany) Sparla, Eric (Germany) Speck, Frank-Olme (Portugal) Speh, Birgit E. M. (USA)

Speicher, Roland (Germany) Speier, Joseph D. (USA) Spengler, Andreas (Germany) Sperber, Wolfram (Germany) Spichtinger, Peter (Germany) Spielberg, Holger (France) Spitkovsky, Ilya M. (USA) Spitzweck, Markus (Germany) Sprekels, Jürgen (Germany) Spresser, Diane M. (USA) Springborn, Boris A. (Germany) Srinivas, Vasudevan (India) Srinivasan, Bhama (USA) Stadje, Wolfgang (Germany) Stafford, Richard M. (USA) Stagnaro, Ezio (Italy) Stagnii, Valerii (Ukraine) Stahlke, Colin (Germany) Stallmann, Friedemann W. (USA) Stamm, Roland (Germany) Stangler, Sieghart S. (Germany) Stanley, Donald (Germany) Stanley, Richard P. (USA) Stark, Harold M. (USA) Staude, Ulrich (Germany) Stawicki, Michael (Germany) Stefânsson, Jôn R. (Iceland) Steger, Tim J. (Italy) Stein, Elias M. (USA) Stein, Michael R. (USA) Stein, Oliver (Germany) Stein, Ralf (Germany) Steinbach, Marc C. (Germany) Steinbart, Enid M. (USA) Steiner, Hans-Georg (Germany) Steiner, Richard J. (United Kingdom) Steinhorn, Charles I. (USA) Stelter, Andreas (Germany) Stempien, Zdzislaw (Poland) Stephan, Holger (Germany) Stephanidis, Nikolaus K. (Greece) Stern, Eugene (USA) Stern, Ronald J. (USA) Sternheimer, Daniel H. (France) Steuding, Jörn (Germany) St. Mary, Donald F. (USA) Stock, Josef K. (Germany) Stocke, A. Britt-Marie (Sweden) Stocks, Douglas R. (USA) Stoer, Josef (Germany)

Stolle, Hermann (Germany) Stoppel, Hannes (Germany) Stork, Frederik (Germany) Storvick, David A. (USA) Stoyan, Yuri Grigoriyevich (Ukraine) Strade, Helmut (Germany) Strano, Rosario (Italy) Strassen, Volker (Germany) Stratigos, Panagiotis D. (USA) Strauch, Matthias (Germany) Strauss, Vladimir (Venezuela) Stray, Arne (Norway) Strazdins, Indulis (Latvia) Streit, Manfred F. G. (Germany) Strey, Gerhard (Germany) Strömberg, Jan-Olov (Norway) Strooker, Jan R. (Netherlands) Stroth, Gernot (Germany) Strugar, Igor (Yugoslavia) Strunkov, Sergey P. (Russian Federation) Struwe, Michael (Switzerland) Strzelecki, Pawel (Poland) Stüben, Hinnerk (Germany) Stühler, Andrea (Germany) Subbotin, Igor Ya. (USA) Suchlandt, Hans E. (Germany) Suciu, Ioan (Romania) Suda, Hiroshi (Japan) Sudan, Madhu (USA) Sugawa, Toshiyuki (Japan) Sugawara, Kazuko (Japan) Suger Cofino, Eduardo (Guatemala) Sujatha, R. (Germany) Sullivan, John M. (USA) Sultangazin, Umirzak (Kazakhstan) Sumi, Hiroki (Japan) Sun, Jialin (P. R. China) Sundberg, Carl (USA) Sunder, V. S. (India) Suresh, Venapally (India) Surkova, Anna (Ukraine) Sustar Rossetti, Ana M. (Argentina) Svendsen, John A. (Norway) Swallow, John (USA) Swanson, Irena (USA) Swart, Henda (South Africa) Swiatek, Grzegorz (USA) Swiatkowski, Jacek (Poland) Sy, Polly W. (Philippines) Symeonidis, Eleutherius E. (Germany)

Symington, Margaret F. (USA) Sytchev, Mikhail A. (Germany) Sytova, Svetlana N. (Belarus) Szabo, Zoltan Istvan (Hungary) Szegedy, Balazs (Hungary) Szemberg, Tomasz (Poland) Sznitman, Alain-Sol (Switzerland) Szolnoki, Dietmar (Germany)

Tachikawa, Atsushi (Japan) Taflin, Erik (France) Taft, Earl J. (USA) Taghavi, Mohsen (Iran) Tai, Yung-sheng (USA) Takahashi, Satoshi (Japan) Takahashi, Sechiko S. (Japan) Takakura, Tatsuru (Japan) Takashima, Keizo (Japan) Takenaka, Shigeo (Japan) Takenouchi, Osamu (Japan) Takeuchi, Kiyoshi (Japan) Talagrand, Michel (France) Taljaard, Monica (South Africa) Tamamura, Akie (Japan) Tamanoi, Hirotaka (USA) Tambieva, Jannet A. (Russian Federation) Tamrazov, Promarz M. (Ukraine) Tanaka, Hisao (Japan) Tanaka, Minoru M. T. (Japan) Tang, Shanjian (P. R. China) Tang, Zizhou (P. R. China) Taniguchi, Yoshiharu (Japan) Tankeev, Serguei G. (Russian Federation) Tao, Terence C. (USA) Tapiz, Chiney (Philippines) Taqqu, Murad S. (USA) Tarasov, Vitaly (Russian Federation) Taraz, Anusch (Germany) Tarkhanov, Nikolai (Germany) Tasaki, Hiroyuki (Japan) Taubes, Clifford Henry (USA) Tavgen, Oleg I. (Belarus) Tchourkin, Andrey V. (Russian

Federation) Tedorov, Kaltscho (Bulgaria) Teicher, Mina (Israel) Terakawa, Hiroyuki (Japan) Terng, Chuu-Lian (USA) Terras, Audrey A. (USA) Teschke, Lothar (Germany) Teschke, Olaf (Germany)

Teuchert, Margitta (Germany) Thalhammer, Mechthild (Austria) Thao, Vo Dang (Vietnam) Thas, Joseph Adolphe François (Belgium) Thomas, Charles B. (United Kingdom) Thomason, Andrew G. (United Kingdom) Thomassen, Carsten (Denmark) Thomsen, Andreas (Germany) Tiba, Dan (Romania) Tienari, Matti (Finland) Tietz, Horst (Germany) Tietze, Walter (Germany) Tijun, Xiao (P. R. China) Tillmann, Heinz-Günther (Germany) Tilouine, Jacques (France) Timmreck, Dagmar (Germany) Timofeeva, Galina A. (Russian

Federation) Timoney, Richard M. (Ireland) Tirado, Eleazar (Venezuela) Tita, Nicolae (Romania) Titzhoff, Frank (Germany) Tkachenko, Natalia Valeryevna (Ukraine) Tkachenko, Vadim A. (Israel) Tkachenko, Viktor Ivanovich (Ukraine) Tkachev, Vladimir G. (Russian

Federation) Tobies, Renate (Germany) Toda, Masahito (Japan) Todorcevic, Stevo B. (Canada) Toffalori, Carlo (Italy) Togari, Yoshio (Japan) Tohge, Kazuya (Japan) Tokizawa, Masamichi (Japan) Tolev, Doytchin (Bulgaria) Toll, Charles H. (USA) Tolok, Alexey (Ukraine) Tolokonnikov, Vadim (Germany) Tomczak-Jaegermann, Nicole (Canada) Tomida, Masamichi (Japan) Tominaga, Yasuo (Japan) Tondeur, Philippe (USA) Ton-That, Tuong (USA) Topalov, Peter (Bulgaria) Törner, Günter (Germany) Totaro, Burt (USA) Totok, Alexander A. (Russian Federation) Toyoshiba, Hiroyoshi (Japan) Tralle, Aleksy (Poland) Trautmann, Günther (Germany)

Trefethen, Lloyd N. (United Kingdom) Trenogin, Vladilen A. (Russian

Federation) Trittler, Matthias (Germany) Trivedi, Vijaylaxmi (India) Troitsky, Evgenij V. (Russian Federation) Tröltzsch, Fredi (Germany) Tronel, Gérard H. (France) Trotman, David J. A. (France) Trott, Michael (USA) Trunk, Carsten (Germany) Tsarkov, Yevgeny (Latvia) Tschinkel, Yuri (USA) Tsevendorj, Ider (Mongolia) Tsirelson, Boris (Israel) Tsou, Sheung Tsun (United Kingdom) Tsuboi, Takashi (Japan) Tsuji, Takeshi (Japan) Tsukui, Yasuyuki (Japan) Tumulka, Roderich A. (Germany) Tuneke, Hilmar R. (Germany) Tung, Chia-Chi (USA) Turaev, Dmitry (Germany) Turo, Jan (Poland) Tuschmann, Wilderich (Germany) Tuzhilin, Alexei A. (Russian Federation) Twardowska, Krystyna (Poland) Tylli, Hans-Olav I. (Finland) Tzafriri, Lior (Israel) Uetake, Yoichi (Poland) Uetz, Marc (Germany) Uhle, Manfred (Germany) Uhlmann, Gunther A. (USA) Ukon, Michihisa (Japan) Ulecia, Teresa (Spain) Ullrich, Peter (Germany) Unger, Luise (Germany) Uraltseva, Nina N. (Russian Federation) Urbach, Paul (Netherlands) Urbonas, Alsimanjas Pranas (Lithuania) Usmanov, Shukhrat (Uzbekistan) Vaccaro, Michelangelo (Italy) Vafa, Cumrun (USA) Väisälä, Jussi (Finland) Valadier, Michel (France) Valla, Giuseppe (Italy) Valtr, Pavel (Czech Republic) van der Geer, Gerard (Netherlands) van der Hoeven, Joris (France) van der Kallen, Wilberd L. (Netherlands) van der Poorten, Alfred J. (Australia) Vanecek, Antonin (Czech Republic) van Gelderen, J. A. (Netherlands) van Lint, Jacobus H. (Netherlands) Vanoli, Marie (France) van Rienen, Ursula (Germany) Varilly, Joseph C. (Costa Rica) Varona, Juan L. (Spain) Varvak, Mark (USA) Vasco, Carlos E. (Colombia) Vasilchuk, Vladimir (Ukraine) Vasil’ev, Alexander Jur’evich (Colombia) Vasiliev, Mikhail (Russian Federation) Vasiliev, Oleg (Russian Federation) Vasilieva, Ekaterina A. (Belarus) Vasilieva, Olga O. (Colombia) Vasilyeva, Tatiana Anatolyevna (Russian

Federation) Veeramani, Palanichamy (India) Velenik, Yvan (Germany) Veloso, Jose Miguel (Brazil) Vélu, Jacques (France) Vemuri, Murali K. (USA) Venema, Gerard A. (USA) Vera Valenzuela, Jaime A. (Chile) Verbitsky, Misha S. (France) Vergara Caffarelli, Giorgio (Italy) Vergne, Michele F. (France) Verrill, Helena A. (Germany) Vershik, Anatoly (Russian Federation) Vershinin, Vladimir V. (Russian

Federation) Vesnin, Andrei Yu. (Russian Federation) Vetter, Udo (Germany) Veys, Willem (Belgium) Viana, Marcelo (Brazil) Vielhaber, Michael J. (Germany) Vignoli, Alfonso (Italy) Viharos, László (Hungary) Villani, Vinicio (Germany) Villella, José A. (Argentina) Vilonen, Kari (USA) Vinberg, Ernest B. (Germany) Vinogradova, Natalya V. (Russian

Federation) Viola, Carlo (Italy) Virag, Balint (USA) Vlad, Carmen D. (USA) Vladimir, Samovol (Russian Federation) Vladimirescu, Mihai M. (Canada)

Vladimirov, V. S. (Russian Federation) Vodop’yanov, Serguei (Russian

Federation) Voevodsky, Vladimir (USA) Vogel, Hans-Jürgen (Germany) Vogelsang, Rudolf (Germany) Vogt, Dietmar (Germany) Vogt, Elmar (Germany) Vogt, Rainer (Germany) Voit, Michael P. (Germany) Volcheck, Emil J. (USA) Volcic, Aljosa (Italy) Volk, Wolfgang (Germany) Vollrath, Hans-Joachim (Germany) Volz, Dietmar (Germany) von Mouche, Pierre (Netherlands) von Schroeders, Nicolai (Germany) von Törne, Christian (Germany) Vortisch, Wilhelm (Germany) Voskresenskii, Valentin E. (Russian

Federation) Voss, Leonard (Germany) Vukovic, Mirjana (Bosnia and

Herzegovina) Waadeland, Haakon S. (Norway) Wacker, Markus (Germany) Wada, Hidekazu (Japan) Wagler, Annegret (Germany) Wagner, Jacques (France) Wagner, Ulrich (Germany) Wagner, Wolfgang (Germany) Wahl, Jonathan M. (USA) Wainger, Stephen (USA) Wakabayashi, Yoshiko (Brazil) Wakimoto, Minoru (Japan) Wakolbinger, Anton (Germany) Walczak, Pawel (Poland) Walczak, Zofia (Poland) Walden, Rainer (Germany) Waldhausen, Maja (Germany) Waldschmidt, Michel (France) Wallace, David A. R. (United Kingdom) Walter, Michael (Germany) Walters, Samuel G. (Canada) Wan, Zhe-Xian (Sweden) Wang, Bai-Ling (Australia) Wang, Feng-Yu (P. R. China) Wang, Hwai-Chiuan (Taiwan) Wang, Lan (USA) Wang, Ya-Guang (P. R. China)

Wang, Zeke (P. R. China) Wängefors, Magnus (Sweden) Wanger, Anke (Germany) Wanka, Gert (Germany) Warmuth, Elke (Germany) Warmuth, Walter (Germany) Wassermann, Gordon (Germany) Watanabe, Hisao (Japan) Watanabe, Shuji (Japan) Watanabe, Toshiro (Japan) Watson, Bill (USA) Watson, Richard O. (Ireland) Waugh, Jason M. (USA) Webber, David (United Kingdom) Weber, Georg (Germany) Wegert, Elias (Germany) Wegmann, Helmut (Germany) Wegner, Bernd (Germany) Wei, Guofang (USA) Weibel, Charles A. (USA) Weigl, Thomas (Germany) Weihmann, Nadja Ilka (Germany) Weikard, Rudi (USA) Weill, Georges G. (USA) Weimann, Lutz (Germany) Weimar-Woods, Evelyn (Germany) Weintraub, Steven H. (USA) Weiser, Martin (Germany) Weisfeld, Morris (USA) Weispfenning, Volker (Germany) Weiss, Jürgen (Germany) Weizsäcker, Heinrich (Germany) Welch, Philip D. (Japan) Welker, Volkmar (Germany) Wellin, Paul R. (USA) Welzl, Emo (USA) Wenk, Carola (Germany) Wenzel, Jörg (Germany) Werbowski, Jaroslaw (Poland) Werner, Dirk (Germany) Werner, Wend (Germany) Wessäly, Roland (Germany) Westerhoff, Natascha (Germany) Westphal, Ursula (Germany) Wieczorek, Tina (Germany) Wiedorn, Corinna (Germany) Wiegand, Sylvia M. (USA) Wiegmann, Klaus Werner (Germany) Wiehe, Martin (Germany) Wieners, Christian (Germany)

Wihstutz, Volker (USA) Wildenhain, Günther (Germany) Wiles, Andrew J. (USA) Wilhelm, Hannes (Germany) Wilke, Marc (Germany) Wilkens, Bettina (Germany) Wilkens, David L. (United Kingdom) Wilkie, Alex James (United Kingdom) Wilkins, David R. (Ireland) Willems, Jan C. (Netherlands) Willerton, Simon (France) Williams, Ruth J. (USA) Wilmanski, Krzysztof (Germany) Wilson, John S. (United Kingdom) Wilson, Pelham M. H. (United Kingdom) Winczewski, Vincent (Germany) Winkelnkemper, Horst E. (USA) Winkler, Jörg (Germany) Winkler, Renate C. (Germany) Winterfeld, Ulrike (Germany) Winterhof, Arne (Germany) Winzer, Olga (Ukraine) Witt, Ingo (Germany) Witte, Dave (USA) Witte, Jürgen (Germany) Wittwer, Janine (USA) Woit, Peter (USA) Wojciechowky, Michal (Poland) Woldar, Andrew J. (USA) Wolf, Joerg (Germany) Wolf, Joseph A. (USA) Wolfart, Jürgen (Germany) Wolfrum, Matthias (Germany) Wong, Siman (USA) Woodin, W. Hugh (USA) Woods, E. James (Germany) Wotzlaw, Lorenz (Germany) Wright, Margaret H. (USA) Wschebor, Mario (Uruguay) Wu, Ching-Tang (Germany) Wu, Jiahong (USA) Wu, Pei Yuan (Taiwan) Wu, Sijue S. (USA) Wu, Siye (Australia) Wu, Xinsheng (P. R. China) Wulff, Claudia (Germany) Wulfsohn, Aubrey (United Kingdom) Wünsch, Volkmar (Germany) Wüstner, Michael (Germany) Wynands, Alexander (Germany)

Xambó-Descamps, Sebastià (Spain)

Xia, Zhihong (USA) Xie, Quanbo (Germany) Ximin, Liu (P. R. China) Xu, Fei (P. R. China) Xu, Xiaoping (P. R. China) Xu, Zhongling (P. R. China)

Yadrenko, Mikhailo (Ukraine)

Yafaev, Dmitri (France)

Yagola, Anatoly G. (Russian Federation) Yagunov, Serge A. (Russian Federation) Yamanaka, Takeharu (Japan) Yamanoshita, Tsuneyo (Japan) Yamashima, Shigeho (Japan) Yamashita, Hajime (Japan) Yan, Min (P. R. China) Yanchuk, Sergiy, V. (Ukraine) Yang, Chung-Chun (P. R. China) Yang, Jae-Hyun (Korea) Yang, Lo (P. R. China) Yang, Tonghai (USA) Yang, Weiqi (P. R. China) Yanusi, Mahmad Yusuf (Tajikistan) Yaroshevsky, Vassily A. (Russian

Federation) Yasuda, Julia Y. (USA) Yasuhiro, Asoo (Japan) Yau, Horng-Tzer (USA) Yebdri, Mustapha (Algeria) Yemelyanova, Inna Sergeevna (Russian

Federation) Yengibarran, Norayi (Georgia) Yevtouchik, Leonik (Russian Federation) Ylinen, Kari E. (Finland) Yoshida, Katsuaki (Japan) Yoshida, Tomoyuki (Japan) Yoshimoto, Akinori (Japan) Yoshino, Masafumi (France) Yost, David (Thailand) Yu, Guoliang (USA) Yu, Jie-Tai (P. R. China) Yu, Joseph M. L. (Canada) Yu, Kunrui (P. R. China) Yu, Pei (Canada) Yu, Rupert W. T. (France) Yugai, Lev (Uzbekistan) Yurko, Vjacheslav A. (Russian Federation) Yushkevich, Alexander A. (USA)

Zacharias, Klaus (Germany)

Zacher, Giovanni G. (Italy) Zagier, Don, B. (Germany) Zaicev, Mikhail V. (Russian Federation) Zaitsev, Dmitri (Germany) Zaitsev, Iuri M. (Brazil) Zakharov, Victor V. (Russian Federation) Zand, Hossein (United Kingdom) Zapatrin, Roman (Russian Federation) Zarbaliev, Sakhavat Mail (Russian

Federation) Zastrow, Andreas (Germany) Zavalishchin, Stanislav T. (Russian

Federation) Zayed, Ahmed I. (USA) Zdravkovska, Smilka (USA) Zegarlinski, Boguslaw (United Kingdom) Zeidler, Eberhard (Germany) Zelazko, Wieslaw (Poland) Zelevinsky, Andrei V. (USA) Zelikin, Mikhail (Russian Federation) Zelman, Mark (USA) Zeman, Martin (Germany) Zemanek, Jaroslav (Poland) Zepf, Günther (Germany) Zerouali, Elhassan (Morocco) Zgliczynski, Piotr (Poland) Zhang, Gong Qing (P. R. China) Zhang, Jian (USA) Zhang, Jinyu (P. R. China) Zhang, Meirong (P. R. China) Zhang, Pu (P. R. China) Zhang, Shou-Wu (USA) Zhang, Shunian (P. R. China) Zhao, Jennifer J. (USA)

Zhao, Yiquiang Q. (Canada) Zhao, Yue (USA) Zharinov, Victor V. (Russian Federation) Zhdanov, Renat Z. (Ukraine) Zheng, Fangyang (USA) Zhu, Jianping (USA) Ziegenbalg, Jochen (Germany) Ziegler, Günter M. (Germany) Ziegler, Martin (Germany) Zielke, Roland J. (Germany) Zillmann, Désirée (Germany) Zimmer, Johannes (Germany) Zimmermann, Bernd (Germany) Zimmermann, Uwe T. (Germany) Zinchenko, Olga A. (Russian Federation) Zink, Wilhelm (Germany) Zipperer, Jörg (Germany) Zitzwitz, Gerhard (Germany) Ziza, Oxana A. (Russian Federation) Zizchenko, Alexei B. (Russian Federation) Zoch, Avery (USA) Zolotykh, Andrej A. (Russian Federation) Zorn, Erhard (Germany) Zouling, Zhou (P. R. China) Zowe, Jochem (Germany) Zschiedrich, Lin (Germany) Zschiegner, Marc (Germany) Zseby, Siegfried (Germany) Zubrilin, Konstantin A. (Russian

Federation) Zuev, Sergey V. (Russian Federation) Zvyagin, Victor G. (Russian Federation) Zwahlen, Bruno (Switzerland) Zygmunt, Marcin J. (Poland)

Participants by Country

(according to their mailing address)

Algeria . . . . . . . . . . . . . . . . . . 5 Argentina . . . . . . . . . . . . . . 15 Armenia . . . . . . . . . . . . . . . . 4 Australia. . . . . . . . . . . . . . . 18 Austria . . . . . . . . . . . . . . . . 17 Azerbaijan . . . . . . . . . . . . . . 1 Bangladesh. . . . . . . . . . . . . . 1 Barbados. . . . . . . . . . . . . . . . 1 Belarus . . . . . . . . . . . . . . . . 12 Belgium . . . . . . . . . . . . . . . . . 9 Bolivia . . . . . . . . . . . . . . . . . . 1 Bosnia and Herzegovina . 2 Brazil . . . . . . . . . . . . . . . . . . 42 Bulgaria . . . . . . . . . . . . . . . . 6 Burundi . . . . . . . . . . . . . . . . . 1 Cameroon . . . . . . . . . . . . . . . 2 Canada . . . . . . . . . . . . . . . . 45 Chile. . . . . . . . . . . . . . . . . . . 11 Colombia. . . . . . . . . . . . . . . . 4 Costa Rica . . . . . . . . . . . . . . 1 Croatia . . . . . . . . . . . . . . . . . 7 Cuba. . . . . . . . . . . . . . . . . . . . 1 Cyprus . . . . . . . . . . . . . . . . . . 3 Czech Republic . . . . . . . . 11 Denmark. . . . . . . . . . . . . . . 11 Egypt . . . . . . . . . . . . . . . . . . . 7 Estonia . . . . . . . . . . . . . . . . 13 Finland . . . . . . . . . . . . . . . . 27 France . . . . . . . . . . . . . . . . 138 Georgia . . . . . . . . . . . . . . . . 13 Germany . . . . . . . . . . . . 1087 Ghana . . . . . . . . . . . . . . . . . . 1 Greece . . . . . . . . . . . . . . . . . 10 Guatemala . . . . . . . . . . . . . . 4 Hungary . . . . . . . . . . . . . . . 25 Iceland . . . . . . . . . . . . . . . . . . 2 India. . . . . . . . . . . . . . . . . . . 35 Indonesia. . . . . . . . . . . . . . . . 1 Iran. . . . . . . . . . . . . . . . . . . . 24 Ireland . . . . . . . . . . . . . . . . . 10 Israel . . . . . . . . . . . . . . . . . . 36 Italy . . . . . . . . . . . . . . . . . . . 65 Ivory Coast . . . . . . . . . . . . . 5 Japan. . . . . . . . . . . . . . . . . 205 Kazakhstan . . . . . . . . . . . . . 7 Kingdom of Saudi-Arabia 5 Korea . . . . . . . . . . . . . . . . . . 10 Kuwait . . . . . . . . . . . . . . . . . . 3 Latvia . . . . . . . . . . . . . . . . . . 6 Lebanon . . . . . . . . . . . . . . . . 1 Lesotho . . . . . . . . . . . . . . . . . 1 Lithuania . . . . . . . . . . . . . . . 6 Luxembourg . . . . . . . . . . . . 2 Macedonia . . . . . . . . . . . . . . 1 Malaysia . . . . . . . . . . . . . . . . 1 Mexico . . . . . . . . . . . . . . . . . 15 Moldova. . . . . . . . . . . . . . . . . 3 Mongolia . . . . . . . . . . . . . . . . 2 Morocco . . . . . . . . . . . . . . . . 9 Mozambique . . . . . . . . . . . . 3 Netherlands . . . . . . . . . . . . 19 New Zealand . . . . . . . . . . . . 5 Nigeria . . . . . . . . . . . . . . . . . . 4 Norway . . . . . . . . . . . . . . . . 18 P. R. China . . . . . . . . . . . . 52 Pakistan . . . . . . . . . . . . . . . . 1 Peru . . . . . . . . . . . . . . . . . . . . 1 Philippines . . . . . . . . . . . . . . 4 Poland . . . . . . . . . . . . . . . . . 74 Portugal . . . . . . . . . . . . . . . 18 Romania . . . . . . . . . . . . . . . 43 Russian Federation …. 202 Senegal . . . . . . . . . . . . . . . . . 1 Singapore . . . . . . . . . . . . . . . 4 Slovakia . . . . . . . . . . . . . . . . . 2 Slovenia . . . . . . . . . . . . . . . . . 1 South Africa . . . . . . . . . . . 17 Spain . . . . . . . . . . . . . . . . . . 44 Sweden . . . . . . . . . . . . . . . . 37 Switzerland . . . . . . . . . . . . 39 Taiwan. . . . . . . . . . . . . . . . . 11 Tajikistan . . . . . . . . . . . . . . . 8 Tanzania . . . . . . . . . . . . . . . . 1 Thailand . . . . . . . . . . . . . . . . 3 Togo . . . . . . . . . . . . . . . . . . . . 1 Tunesia . . . . . . . . . . . . . . . . . 6 Turkey . . . . . . . . . . . . . . . . . . 8 Turkmenistan . . . . . . . . . . . 1 USA . . . . . . . . . . . . . . . . . . 500 Ukraine . . . . . . . . . . . . . . . . 75 United Arab Emirates … 1 United Kingdom . . . . . . . 88 Uruguay . . . . . . . . . . . . . . . . 5 Uzbekistan . . . . . . . . . . . . . . 6 Venezuela . . . . . . . . . . . . . . . 5 Vietnam . . . . . . . . . . . . . . . . 5 Yugoslavia . . . . . . . . . . . . . 21 Zimbabwe . . . . . . . . . . . . . . . 2